### v1

# Prompts to Proxies: Emulating Human Preferences via a Compact LLM Ensemble

Abstract
Large language models (LLMs) have demonstrated promise in emulating human-like responses across a wide range of tasks. In this paper, we propose a novel alignment framework that treats LLMs as agent proxies for human survey respondents, affording a cost-effective and steerable solution to two pressing challenges in the social sciences: the rising cost of survey deployment and the growing demographic imbalance in survey response data. Drawing inspiration from the theory of revealed preference, we formulate alignment as a two-stage problem: constructing diverse agent personas called endowments that simulate plausible respondent profiles, and selecting a representative subset to approximate a ground-truth population based on observed data. To implement the paradigm, we introduce P2P, a system that steers LLM agents toward representative behavioral patterns using structured prompt engineering, entropy-based sampling, and regression-based selection. Unlike personalization-heavy approaches, our alignment approach is demographic-agnostic and relies only on aggregate survey results, offering better generalizability and parsimony. Beyond improving data efficiency in social science research, our framework offers a testbed for studying the operationalization of pluralistic alignment. We demonstrate the efficacy of our approach on real-world opinion survey datasets, showing that our aligned agent populations can reproduce aggregate response patterns with high fidelity and exhibit substantial response diversity, even without demographic conditioning.
Keywords Preference modeling  $\cdot$ Value alignment  $\cdot$ Evaluation
1 Introduction
Pollsters are facing mounting challenges in the contemporary landscape, as public willingness to participate in surveys declines across regions and over time. In the United States, telephone response rates have dropped from 36% in the late 1990s to under 6% today Kennedy (2019), while major federal surveys such as the Current Population Survey (CPS) have experienced a 20-point drop since 2015 U.S. Bureau of Labor Statistics (2025). Likewise, across Europe Jabkowski and Cichocki (2025), response rates for national surveys in 36 European countries have fallen 0.8% per year on average between 1999 and 2018. This erosion of participation undermines statistical power and representativeness, particularly because the decline is not uniform across demographic groups. For instance, recent election polling failures have been traced to differential nonresponse: Republicans and independents are increasingly less likely to answer than Democrats Clinton et al. (2022); Parshall (2024). In the social sciences, this is a classic case of sample selection bias, which, if left unchecked, compromises both predictive accuracy and external validity of the survey-based inference. At the same time, the exorbitant cost of conducting large-scale surveys, due to respondent recruitment, compensation, and extensive piloting, places constraints on experimentation and iteration Kennedy (2015). These challenges highlight a pressing need for scalable, flexible, and theoretically grounded alternatives for modeling human preferences.
The epistemic challenge of capturing human preferences, long studied in the social sciences, now finds a structural parallel in large language model (LLM) alignment, where researchers aim to steer model behavior in line with human values and beliefs. Standard approaches, such as supervised fine-tuning (SFT) Tan et al. (2024) and reinforcement learning from human feedback (RLHF) Bai et al. (2022); Christiano et al. (2017), train models to reflect normative ideals of helpfulness, harmlessness, or truthfulness. While effective for building general-purpose assistants, these techniques are inherently monolithic: they produce a single dominant behavioral profile, making them ill-suited for modeling pluralistic human populations. Scaling them to reflect multiple value systems requires retraining or fine-tuning for each group, rendering the approach costly, brittle, and inflexible. Consequently, these pipelines struggle with tasks that demand population-level diversity, such as simulating public opinion or generating group-sensitive predictions.
To address this rigidity, alignment researchers have increasingly turned to post-training conditioning techniques that guide model behavior at inference time. A prominent example is the use of structured prompts, variously termed endowments Horton (2023), personas Castricato et al. (2025), or Silicon Samples Argyle et al. (2023), that encode sociodemographic attributes or behavioral dispositions. Prompted this way, LLMs can emulate individual-level responses with considerable fidelity Bisbee et al. (2024), making them attractive for constructing synthetic populations to pilot surveys or test hypotheses. However, this approach remains limited in key respects. First, it depends heavily on human-curated prompts or ground-truth demographic profiles, which may be unavailable or incomplete. Second, it often produces synthetic responses that lack the natural diversity of real populations: LLMs exhibit central-category bias Wang et al. (2024) and generate insufficient inter-agent variation Bisbee et al. (2024). Thus, while endowments reduce the need for retraining, they increase reliance on demographic templates and fall short of capturing true representational diversity.
These limitations reveal a common assumption in alignment: that behavioral fidelity hinges on demographic accuracy. We challenge this view and propose a new theoretical framework grounded in preference reconstruction. Rather than aligning each synthetic agent to a ground-truth demographic profile, we argue that effective emulation requires only the recovery of the aggregate preference structure of the target population. Drawing on revealed preference theory in economics, we treat each agent as a proxy decison-maker whose responses expose latent beliefs or utilities. Crucially, we do not seek to recover the true basis of individual human preferences. Instead, it suffices to identify a functional basis in the latent preference space, a diverse set of proxy agents that span the opinion manifold, and determine appropriate aggregation weights over them. This abstraction is reflected in human societies: demographically distinct populations may exhibit similar collective attitudes, even if their constituent beliefs diverge. Our approach thus enables a more data- and alignment-efficient route to survey emulation, moving beyond identity-conditioned generation toward flexible distributional modeling.
This perspective also clarifies and concretizes the goals of pluralistic alignment Sorensen et al. (2024), which advocates representing a spectrum of human values rather than optimizing for a singular objective. While compelling in principle, the pluralistic alignment literature has remained largely agnostic about how to measure pluralism or evaluate pluralistic systems. Our framework addresses this gap by proposing concrete criteria for measuring variation and fidelity, along with a mechanism to steer agent distributions efficiently under resource constraints so that the aggregate distribution of answers across the agent ensemble matches that of a given target population. In doing so, we operationalize distributional pluralistic alignment through a lens grounded in economic theory and suited for practical survey applications.
Our contributions are as follows:
  • Theoretical foundation. We formalize preference alignment as a two-stage process: constructing a functional basis of proxy agents and determining aggregation weights, drawing on revealed preference theory to recover population-level preferences.
  • Modular alignment system. We implement this theory via a modular system, Prompts to Proxies (P2P), that combines active endowment generation (to promote response diversity) with regression-based aggregation (to estimate preferences without demographic alignment).
  • Pluralistic alignment in practice. We provide the first operational framework for pluralistic alignment, introducing metrics for diversity and fidelity and propose scalable strategies to balance representational coverage and resource efficiency.
  • Bridge to empirical social science. Our framework enables social scientists to simulate and test preference theories by controlling agent endowments and tracking their effects on response distributions. It also supports survey design and can mitigate nonresponse bias, offering a cost-effective complement to traditional data collection.
  • Testbed for alignment research. The modular design provides a controlled environment for evaluating prompting strategies, enabling alignment researchers to explore steerability and value-sensitive modeling at scale.
2 Modeling Preferences via Functional Bases
Preference modeling under conventional lenses
Human preferences are latent, complex, and often inconsistent, posing a persistent challenge for formal modeling. Economists and social scientists have approached this challenge through multiple modeling frameworks. One prominent approach models humans as rational agents who maximize a stable utility function. While elegant in theory and mathematically tractable, these models often struggle to explain empirical anomalies in observed behavior Kahneman and Tversky (1979); Thaler (1999); Kahneman and Tversky (2000). A complementary framework, revealed preference theory Samuelson (1938); Mas-Colell et al. (1995), shifts the focus from explicit utility functions to observable choices. It asks whether a pattern of choices can be rationalized as if generated by a utility-maximizing agent without assuming a specific functional form. This idea has proven especially valuable in empirical settings, offering a way to assess whether behavior is consistent with stable preferences directly from observed data.
We take inspiration from the revealed preference perspective—not in its full rationality axioms, but in its core idea: preferences, though hidden, can be inferred from behaviors, such as responses to survey questions. This principle also underlies modern alignment pipelines such as SFT and RLHF, where models are steered toward human-preferred outputs using behavioral signals in the form of labeled data or feedback on generated responses.
Our framework
To formalize our approach, we zoom out from individual preferences and return to the level of survey analysis. Surveys are not designed to recover each respondent’s utility function; rather, they aim to capture aggregate population preferences, often through summary statistics such as response distributions or marginal means. This shift in perspective from individual to population invites a functional view (Figure 1): we can represent the population preference as a linear combination of individual preference vectors in a latent preference space. Responses to survey questions can then be viewed as mappings from this latent space to an observable response space. Crucially, this mapping is generally many-to-one: different underlying preferences may yield identical observed responses. Yet, as the number and diversity of responses increase, the expressive capacity of the response space improves. In the limit, a sufficiently rich battery of responses enables an increasingly precise identification of latent preferences. This perspective motivates our framework: Instead of recovering the ground-truth preference of each individual, we can construct a functional basis, a diverse set of proxy agents whose preferences span the latent space. By observing their responses and estimating the appropriate aggregation weights, which can be positive or negative, we can reconstruct the aggregate preference structure of the population.
Figure 1: Top: Finding a functional basis to recover latent ground-truth preference. Bottom: Identification of latent preference through multiple observations.
Alignment as a two-stage problem
Based on our theoretical framework, we can cast alignment as a two-stage problem. The first stage involves constructing a broad and expressive candidate basis of proxy agents, each representing a distinct vector in the latent preference space. For the basis to be effective, it must capture meaningful variation across the population. This objective aligns closely with recent calls for pluralistic alignment Sorensen et al. (2024), which stresses the importance of representing a spectrum of human values rather than converging on a single ideal. However, we do not have direct access to individual preferences, and the prompt space is far too vast to explore in an unstructured manner. To address this, we adopt a structured prompting strategy based on attributes—interpretable factors whose variation is likely to influence preference expression. These attributes may be informed from theory (e.g., ideology, values, levels of altruism) or content (e.g., topical relevance). By combining and varying these attributes systematically, we generate a set of endowment-conditioned agents that span different positions in the preference space.
The second stage concentrates on designing an aggregation mechanism to reconstruct the population preference from proxy agents. We frame this as a supervised learning problem: given aggregate responses for a subset of survey questions, we estimate a set of aggregate weights over the proxy agents based on their individual response patterns. These weights are then used to predict responses on held-out questions, enabling out-of-sample evaluation. This setup allows us to examine whether the weighted agentic ensemble indeed captures the target population preference, rather than overfitting to the training set. We adopt regression-based methods for this purpose, as they align naturally with our assumption that population preferences emerge as linear combinations of individual vectors in the latent space. We describe implementation details, including attribute construction, endowment generation, and constrained regression methods, in the following section.
3 Prompts to Proxies (P2P)
To implement our preference reconstruction strategy in the survey context, we engineer a modular alignment system called Prompts to Proxies (P2P) (Figure 2). The system consists of two core components: (1) active endowment generation, powered by a dynamic attribute bank, and (2) regression-based aggregation, where constrained variable section algorithms form the backbone. They concretize the two stages of alignment introduced in the previous section and enable the generation of a compact set of opinionated LLM agents to emulate the results of observed population-level surveys.
Figure 2: System Overview of the Prompts to Proxies (P2P) Framework. P2P operationalizes preference alignment as a two-stage process inspired by revealed preference theory. In Stage 1, active endowment generation uses a structured attribute bank and entropy-guided adaptive sampling strategies to construct diverse agent personas, called endowments. In Stage 2, regression-based aggregation assigns weights to reconstruct population-level preference patterns. An external validity check accesses how well the weighted ensemble predicts aggregate responses on held-out questions, assessing out-of-sample generalization.
3.1 Survey Splitting and Evaluation Protocol
To support the alignment procedure, we partition the full set of survey questions, denoted $\mathcal{Q}$, into three disjoint subsets: $\mathcal{Q}_{\mathrm{train}}$, $\mathcal{Q}_{\mathrm{valid}}$, and $\mathcal{Q}_{\mathrm{test}}$, corresponding to the training, validation, and test sets, respectively. Let $N=|\mathcal{Q}|$ denote the total number of questions, and $N_{\mathrm{train}}$, $N_{\mathrm{valid}}$, and $N_{\mathrm{test}}$ the cardinalities of the three subsets. P2P allows users to set the split sets and ratio flexibly.
For the alignment system, $\mathcal{Q}_{\mathrm{train}}$ is used for active endowment generation. $\mathcal{Q}_{\mathrm{train}}$ and $\mathcal{Q}_{\mathrm{valid}}$ are used in regression-based aggregation. We reserve $\mathcal{Q}_{\mathrm{test}}$ as an ultimate test of the predictive accuracy of the selected agent ensemble, ensuring that the reconstructed preference has a decent degree of generalizability. For regression analysis, categorical questions are binarized into question-option pairs.
3.2 Attributes as Control Handles
As a crucial part of the active endowment generation logic, we use attributes to shape model preferences. This approach, though counter-intuitive at first glance, is in fact a deliberate reverse engineering design choice. In humans, attributes such as ideology, dispositions, or values are better understood as reflections or downstream manifestations of latent preferences. However, language models do not possess preferences in the human sense. As next-token prediction machines, their behavior is shaped by statistical associations learned from human text corpora. This distinction allows us to invert the causal direction: rather than treating attributes as reflections of internal preference states, we tap into the pattern matching nature of LLMs and use attributes as control handles to steer model behavior.
To ensure that attributes effectively steer model preferences, we design a preset attribute bank that draws inspiration from theories across diverse social science disciplines, such as economics, political science, management science, and psychology. The bank is organized hierarchically: individual attributes are grouped into modes, each representing a coherent subject area (e.g., values and beliefs, cognitive biases, or political ideology). These modes are further clustered into higher-level templates, which reflect different design purposes or theoretical orientations.
We define three primary templates. The core template contains general demographic descriptors and serves as the default baseline. The thematic template includes topic-specific modes such as economics, politics, culture and history. The theoretical template groups attributes derived from canonical frameworks such as Maslow’s hierarchy of needs Maslow (1943) and the Big Five personality traits John and Srivastava (1999). A condensed summary of these templates, representative modes, and attributes is shown in Table 1. These templates and modes are not fixed: they serve as a structured starting point but are fully extensible. Users can define new templates, introduce domain-specific modes, or modify existing attribute sets to suit different experimental contexts. By sampling and varying attributes from this flexible bank, we create prompts that instantiate distinct proxy agents occupying different regions of the latent preference space.
Template Representative Modes and Attributes
Core core: gender, race, ethnicity, education, income level,
Thematic economics: occupation, price sensitivity, wealth/debt,
politics: policy preferences, civic engagement, ideological orientation,
behavioral_biases: status quo bias, loss aversion, overconfidence bias,
Theoretical maslow: safety needs, self-actualization, love and belonging needs,
big_five: conscientiousness, neuroticism, extraversion,
Table 1: Structured view of the attribute bank, with representative modes and attributes grouped by template. Users can extend the bank with extra templates, modes, or attributes.
While the attribute bank provides structured coverage, we also enable dynamic attribute discovery. We introduce two additional freeform templates that are specific to the survey under analysis. The survey template derives attributes by prompting a designated LLM, called attribute learner, to identify relevant decision-making factors based on the overall survey topic covered by the training set of questions, $\mathcal{Q}_{\mathrm{train}}$. The question template goes further by asking the learner to derive attributes tailored to a specific survey question. In both cases, the inferred attributes are appended to the attribute bank only for the duration of the survey run. This hybrid strategy strikes a balance between theory-driven and data-driven endowment generations, and enables human-AI co-discovery of salient preference descriptors.
With the attribute bank in place, we build an endowment generator using a dedicated LLM termed endowment model that takes a set of attributes as input and produces diverse agent profiles by varying the instantiation of each selected attribute.
3.3 Active Endowment Generation
To construct a diverse preference basis without exhaustively enumerating attribute combinations, we create an active endowment generation procedure to operate under resource constraints. It proceeds in iterative steps. At each step, new endowments are sampled, instantiated as agents to elicit responses to survey questions, and evaluated based on their contribution to response diversity. The remainder of this section describes the key components of this procedure, including our variability scoring strategy, adaptive sampling steps, and entropy-based patching mechanisms.
Tracking response diversity with variability score
As established in Section 2, observed responses can serve as behavioral manifestations for latent preference vectors. Through this link, response variability serves as an initial diagnostic for assessing diversity in the underlying preference space. To effectively measure response variability, we introduce the notion of question entropy:
$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}$ (1)
where $H_{i}(\mathcal{A})$ denotes the normalized entropy of question $i$ based on responses from a group of agents $\mathcal{A}$, $K_{i}$ is the number of unique response options $\{1,\dots,K_{i}\}$ for question $i$, and $p_{ik}$ is the empirical proportion of responses selecting option $k$. The denominator normalizes for scale, ensuring that entropy is comparable across questions with different numbers of choices. A high question entropy indicates greater response diversity for that question. By varying the group $\mathcal{A}$, e.g., restricting it to agents sampled from a particular mode, we can compute question entropy at the group level and compare the diversity contributes across modes.
To turn question entropy into a useful metric for adaptive sampling, we introduce variability score for each mode. Formally, it is defined as
$V(\mathcal{A}_{\mathrm{mode}})=\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\frac{H_{i}(\mathcal{A}_{\mathrm{mode}})}{H_{i}(\mathcal{A}_{\mathrm{total}})}.$ (2)
where $H_{i}(\mathcal{A}_{\mathrm{mode}})$ denotes the entropy of responses to question $i$ based on the cumulative set of agents from a given mode up to the current generation step, and $H_{i}(\mathcal{A}_{\mathrm{total}})$ is the corresponding question entropy across all agents generated so far. Computed this way, the variability score favors modes that are capable of generating endowments which elicit diverse responses to questions that other modes typically fail to diversify. The intuition is to maximize the informational utility of each observation under budget constraints. When most agents agree on a question, a divergent response signals a valuable distinction in the underlying preference space. In P2P, we task a dedicated class object, Tracker, to compute the variability score for each mode after each generation round. The scores are then passed through a softmax function to produce a probabilty distribution over modes, which is used to guide endowment sampling in the next generation step.
Adaptive sampling and update steps
In P2P, endowment generation takes several steps. At the outset of the generation, an attribute learner is invoked to infer attributes from the survey based on the training questions, yielding the survey template on top of the existing templates in the attribute bank. The system then enters an initial sampling stage, drawing an equal number of endowments from each mode. The endowments are then used to instantiate agents to elicit responses to training questions, $\mathcal{Q}_{\mathrm{train}}$. The collective responses are then passed to Tracker for the computation of the variability scores and the sampling probabilities for the next update step. At the next update step, endowment budget allocation follows the multinomial distribution. A higher probability exploits the top-performing mode, while randomness allows exploration of its under-performing peers.
Patching low-entropy questions
In addition to the existing trade-off between exploitation and exploration, we adopt an additional boosting strategy, which we call question patching. At the end of each update step, as the Tracker computes the variability scores for each existing mode, it also keeps track of the questions with the lowest entropy values. When question patching is activated, the attribute learner will be called in the next generation step to exclusively infer attributes on each of the lowest-$n$ entropy questions, where $n$ is user-defined. The split of the endowment budget between the active generation and boosting generation is governed by a user-specified ratio.
Figure 3: Entropy change across update steps for each question during active endowment generation for ATP W42. Each column represents a question. Green indicates an increase in entropy from the previous step, while red indicates a decline.
To avoid wasting resources on persistently low-entropy and to promote cross-pollination between modes, we additionally introduce a mixed mode strategy. It is triggered after a question appears $n$ times in the lowest-$k$ entropy list, where $n$ is a user-defined threshold. When in effect, the strategy relocates the assigned endowment budge to a new mixed mode, which combines attributes from the question with those of the current top-performing mode. This encourages diversity injection from high-performing modes while maintaining a question-specific focus.
The adaptive sampling cycle continues until the generation budget $N_{A}$ (an upper bound on the maximum number of agents allowed) is exhausted. By iteratively targeting low-diversity questions and expanding the attribute space, active endowment generation progressively increases response diversity while preserving coverage over previously explored regions of the preference space.
3.4 Regression-Based Aggregation
With the candidate proxy basis assembled, P2P enters the second stage: Regression-based Aggregation. The goal at this stage is to find a functional basis from the candidate vectors to reconstruct the observed preference. Depending on the endowment budget and the total number of trainable questions $\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}$, we could end up in a situation where there are more variables than observations. Consequently, a variable selection algorithm is vital. Besides, even with $N_{\mathrm{train}}+N_{\mathrm{valid}}>>N_{A}$, variable selection allows us to retain a parsimonious agent ensemble, saving inference costs on further simulations. To conduct variable selection, we consider regression methods with an added L1 penalization term. We can use either hold-out validation or cross-validation (CV) for hyperparameter tuning. The objective of the learning problem can be stated as
$\min_{\mathbf{w}}\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\sum_{k=1}^{K_{i}}\left(r_{ik}-\sum_{j=1}^{N_{A}}w_{j}d_{j,ik}\right)^{2}+\alpha\Omega(\mathbf{w}),$ (3)
$\mathrm{s.t.}\quad\sum_{k=1}^{K_{i}}\sum_{j=1}^{N_{A}}w_{j}d_{j,ik}=1\quad\forall i\in\mathcal{Q}_{\mathrm{train}},$ (4)
where $\mathbf{w}\in\mathbb{R}^{N_{A}}$ is the weight vector for the agents, $r_{ik}$ denotes the ground-truth proportion of responses selecting option $k$ for question $i$, $d_{j,ik}$ is a dummy response variable equal to $1$ if agent $j$ selects option $k$ and 0 otherwise, and $\alpha$ is the hyperparameter for the penalization. When $\Omega$ consists of only a L1 norm, the formulation amounts to constrained lasso. When $\Omega$ subsumes an additional L2 norm, it becomes constrained elastic net. Our simulation studies indicates that constrained lasso faithfully recovers the ground-truth agents from a large candidate pool under an adequate number of observations (Appendix A.1), and it is potent in selecting a proxy ensemble to recover ground truth even when the number of proxy agents is small (Appendix A.2). Both results corroborate using it as a selection backbone. Our simulation study also investigates the relationship between entropy and prediction fit—we leave the details to Appendix A.3 for interested readers.
4 Taking P2P to the Wild
As a first empirical test of the alignment framework, we deploy P2P on real-world social survey data. Specifically, we use the American Trends Panel (ATP), a nationally representative U.S. survey conducted by Pew Research Center in discrete waves. In this demonstration, we choose Wave 42, which probes trust in science by asking respondents about scientific competence, integrity, and public accountability. The wave includes responses from approximately 4464 U.S. adults Pew Research Center (2019) and the cleaned dataset contains 128 questions.
For the experiment, we split the questions according to a $7:1.5:1.5$ ratio into the training, validation and test sets. For the endowment generation stage, we set the endowment budget to be $300$, with $10$ endowments generated for each mode during initial sampling. Subsequently, the rest of the endowments are generated across 5 update steps. In each step, we allocate 75% of the endowment budget to question patching, where the lowest 3 questions are considered. We set the trigger threshold of the mixed mode strategy at 1, so that the budget is reallocated if a question enters the lowest-3 list the second time. At the end of the generation, 5 questions are patched, with additionally 5 mixed modes created.
Endowment generation improves entropy
Figure 3 displays the entropy change for each question across the update steps. Evidently, a significant number of questions experience a noticeable rise in entropy at the end of the first update step. From the second step onward, mixed mode enters the generation loop. Marginal gains in entropy are observed for the majority of questions, whereas some experience mild entropy drops. Appendix B provides question entropy trajectories in levels. The three questions that exhibit the largest drops in entropy are among the highest-entropy questions at the initial sampling stage. The evolution elucidates the tradeoff made by the active endowment generator encoded in the variability scores.
For the selection of the agent ensemble, we fit both a constrained lasso and a constrained elastic net, using CV to select the penalization parameters. Consequently, lasso selects 53 out of the 300 synthetic agents to form the agent ensemble, while elastic net selects 43. The training and testing performances are similar between the two methods.
Figure 4: Prompts to Proxies: Emulation Results for ATP W42. Top: Train (left) and test (right) prediction performances for the binarized responses. Bottom: Human and agent (pre- and post- alignment) aggregated responses for the sample test question RQ4_F1Ae_W42.
Prediction indicates feasibility
Figure 4 shows the prediction performances of the agent ensemble selected with constrained lasso. The train-time root mean squared error (RMSE) is 0.08, with test RMSE slightly higher at 0.09, indicating adequate generalization ability of the functional basis. The bottom panel of Figure 4 presents a snapshot for the aggregated responses for a test set question, comparing the human ground truth with agentic emulations before and after regression-based aggregation. Overall, the results are encouraging, especially considering that P2P constructs the endowments based solely on preset and learned attributes and is agnostic on the ground-truth demographic data.
The panoramic view and future directions
In Appendix B, we provide a preliminary panel study, applying P2P to 14 ATP waves, each repeated 3 times. Performance variation is relatively low within each wave but considerably higher across the panel. In 8 out of 14 waves, P2P consistently achieves a test MSE below 0.015 across all repeats. Four waves exhibit test MSEs fluctuating around 0.015, while the remaining two fall in the range of 0.020 to 0.025. While our simulation studies suggest that entropy coverage, the number of observations and the size of the proxy pool all influence emulation performance, the preliminary panel results signal intriguing directions for future study. Understanding the sources of variation in empirical performance—particularly in relation to both the similarities and the distinctions from simulation results—offers a promising lens on the dynamics of human–AI alignment. These insights can also inform refinements to P2P ’s modular architecture and motivate extensions of its functionality.
P2P as a means to an important end
At a broader level, our framework operationalizes pluralistic alignment Sorensen et al. (2024). Its structured use of attribute-based control enables the representation of diverse value systems, while its regression-based weighting mechanism supports distributional pluralism without enforcing demographic fidelity. Crucially, P2P demonstrates that steerability and pluralism are not separate challenges but deeply intertwined: the ability to steer model behavior through promptable attributes is what enables meaningful diversity in aggregate preference reconstruction.
Although developed in the context of survey emulation, P2P generalizes naturally beyond surveys. The underlying preference reconstruction logic can be applied to other forms of observational data by discretizing continuous variables and binarizing categorical ones. Moreover, while this paper focuses on modeling aggregate population preferences, the framework could be adapted for individual-level alignment—for example, by learning proxy ensembles that emulate preferences of a single person across varied contexts.
Beyond its methodological contributions, P2P affords a modular platform for empirical alignment. Our work is among the first to streamline the use of large-scale social survey data—not as textual corpora for language model pretraining, but as a structured, theory-informed substrate for steering models toward specific populations, cultural milieus, and value contexts. It supports systematic evaluation of prompt engineering techniques, attribute design strategies and sampling heuristics. Researchers can test how different endowment structures affect response diversity, and how entropy metrics reflect or obscure latent preference variation. These tools offer a bridge between high-level alignment goals and grounded behavioral evaluation.
Finally, P2P offers a testbed for social scientists to probe and operationalize preference theories. By investigating how assumptions about cognitive style, value orientation, or belief structure affect response distributions and mode-level entropy, researchers can explore the explanatory power to theoretical constructs in a controlled but flexible setting. In this sense, the system fulfills the spirit of its name: P2P connects peer disciplines, translates prospects into products, carries ideas from pseudocode to paper, and moves pluralistic alignment toward practical needs.
5 Conclusion
This work establishes a theoretical foundation for viewing preference alignment as a problem of preference reconstruction. Inspired by revealed preference theory, our approach formalizes alignment as a two-stage process: first, constructing a functional basis of proxy agents using attribute-guided prompting; and then, recovering aggregation weights through supervised learning. In doing so, our work offers the first tractable implementation of pluralistic alignment, along with quantitative metrics to assess both diversity and fidelity.
Beyond its theoretical contributions, this framework offers a practical platform for interdisciplinary research. In the social sciences, it can support survey design, question testing, and nonresponse mitigation, serving as a cost-effective complement to traditional data collection. For researchers in LLM alignment, the modularity of our system provides a controlled setting to evaluate prompt engineering strategies and their influence on agent diversity and downstream performance. More broadly, it lays the groundwork for agentic systems that enable value-sensitive modeling across disciplines. Notably, P2P repositions social survey data not as training material, but as a guiding signal for culturally grounded model alignment.
Appendices
Appendix A Simulation Studies
In order to gauge the validity of the proposed preference model and the factors affecting alignment performance, we conduct three simulation studies where we have the privilege of ground-truth knowledge.
Simulation Design
For the simulation studies, we use the American Trends Panel (ATP) Wave 42 (W42) as the survey source, covering the topic of public trust in science and scientists. We run the active endowment generation on the survey to get 300 persona endowments, and partition the survey questions into training, validation and test sets by the 0.7:0.15:0.15 ratio. We subsequently simulate the responses to the survey questions using endowed GPT-4o agents as the synthetic respondents. The ATP W42 dataset has in total 129 questions, which after binarization leads to 517 observations—85% of these observations form the train-and-validation (trainval) set and the rest 15% is used to test the predictive accuracy of the selected agent ensemble.
Construction of the Ground Truth Data
To construct the ground truth data, we sample a fraction of the endowments as the ground truth and randomly assign weights to these endowments, which represent the proportions of the associated personae in the ground truth population. The ground truth data—percentages of individuals choosing difference options for a given question—are computed using weighted averages of responses provided by the ground-truth agents. The unselected endowments are used to construct the pool of proxy agents.
Regression Method
For the simulation studies, we choose constrained lasso as our main regression alignment method for investigation. As a secondary choice, constrained elastic net is an extension of constrained lasso that trades off between the L1 and L2 penalty terms. It is expected to have more stable performance when the features (agent responses) are highly correlated. Our initial simulations indicate that constrained elastic net demonstrates near identical performance to constrained Lasso, partly because the latter is a special case of the former and we are using cross-validation to select the best penalization parameters. Therefore, for the sake of building intuitions, we focus on constrained lasso in the simulations and expect the results to be largely indicative of constrained elastic net’s performance.
A.1 Recovering Ground Truth Agents with Constrained Lasso
As a first test of constrained lasso as a variable selection algorithm, we aim to evaluate its ability to recover ground-truth agents from a broader candidate pool. To render the selection task nontrivial, we set the prevalence of the ground-truth agents to 0.3—out of 10 selectable agents, only 3 contribute to the construction of the ground-truth signal. As the proxy agents dominate the agent pool, an adequate number of observations becomes essential for Lasso to reliably identify the true contributors. As a key metric, we define observation-to-agent ratio (OAR) to be the ratio of the number of training and validation observations to the number of selectable agents. By default, the OAR for our simulation baseline is 1.203.
To develop a panoramic view of constrained lasso’s ability in ground truth recovery, we conduct multiple simulation rounds under varying OARs. We generate different OAR variations using two complementary subsampling strategies:
  • Subsampling endowments. To increase the OAR, we adopt a subsampling strategy on endowments, where we subsample from the pool of selectable endowments, varying the total number of candidate agents while keeping the proportion of ground-truth agents fixed at 30%. The ground-truth weights are renormalized to sum to 1. We vary the subsample fraction from 0.5 to 1 in 20 equally spaced steps.
  • Subsampling questions. To decrease the OAR, we subsample survey questions while keeping the original train-validation-test split and ratio unchanged. This mimics a setting with reduced behavioral signal for inference. As in the endowment-based approach, we use 20 subsample fractions ranging from 0.5 to 1.
We repeat each subsampling procedure 10 times to construct the final dataset used in our simulation study. Figure 5 shows the trends of the mean squared error of the test set and the precision—that is, the proportion of selected agents who possess ground-truth endowments—across different OARs. To further illustrate the dynamics of selection, Figure 6 presents three representative snapshots of simulation runs, corresponding to low, medium, and high OAR settings.
As the figures indicate, constrained lasso’s ability to recover ground-truth agents from the candidate pool improves with a higher observation-to-agent ratio. This is intuitive as more observations offer a richer signal of both individual-level idiosyncrasies and aggregate preferences. With abundant behavioral cues, the selection algorithm is better equipped to distinguish true contributors from correlated siblings. In contrast, when observations are sparse, Lasso tends to relinquish its sharpness as a selection tool and instead defaults to identifying a functional basis that best explains the limited data. The mild decline in predictive accuracy under low OAR regimes suggests that this functional basis still approximates the aggregate preference reasonably well, albeit imperfectly.
Figure 5: Mean Squared Error (Test Set) and Precision by Different Observation-to-Agent Ratio (OAR) for Constrained Lasso. Each simulation is repeated 10 times with different random seeds. Subsampling of endowments is done with fixed ground-truth-to-proxy ratio. Subsampling of responses is done with fixed train-valid-test split and ratio.
Figure 6: Snapshots of Weight-vs-Coefficient Comparison and Prediction Performance. Each row represents a simulation run. The red bars on the left panel represent ground truth agent weights used to generate the aggregate data and the blue bars Lasso coefficients. The observation-to-agent ratios for the top, middle and bottom panels are: 0.617, 1.203, 2.407. Lasso’s ability to recover the ground truth agents from the agent pool degrades with lower observation-to-agent ratio, but its predictive accuracy only drops mildly.
A.2 Emulating Ground Truth Responses with Proxy Agents
While the first simulation exercise has readily allured to constrained lasso’s ability to find a functional basis to explain the observed data, in this second simulation study we extend the investigation by removing the ground-truth agents from the selectable agent pool. Under this setting, constrained lasso needs to fully rely on proxy agents to recover the ground-truth patterns. We analyze two distinct scenarios:
  • Subsampling proxies. We fix the total number of observations and reduce the proxy pool by retaining only a fraction as selectable agents. We sweep across the fractions from 0.05 to 1 in 20 equal spaced steps.
  • Subsampling training and validation questions. We fix the proxy pool but retain a fraction of training and validation observations. Similar to subsampling proxies, we construct 20 simulation rounds gradually raising the retained fraction from 0.05 to 1.
Like in Simulation Study 1, we repeat each simulation 10 times using different random seeds to form the final results. In Figure 7, the left panel displays the mean squared errors (MSE) and coefficients of determination ($R^{2}$) for the subsampling proxies scenario, while the right panel showcases the associated metrics for the subsampling training and validation questions scenario.
Figure 7: Left: Mean Squared Errors and Coefficients of Determination ($R^{2}$) by Number of Proxy Agents. Right: Mean Squared Errors and Coefficients of Determination ($R^{2}$) by Number of Observation. Each simulation is repeated 10 times with different random seeds.
Notably, constrained lasso exhibits robust performance across all simulated conditions. Even with only 10 selectable proxy agents, it achieves a mean test MSE of $0.0067$, yielding an $R^{2}$ of $0.93$. Predictive accuracy improves as we increase the number of selectable proxy agents, plateauing at $0.99$. In comparison, decreasing the number of observations has a lesser effect on predictive performance: in the recorded run with the lowest number of observations at 13, constrained lasso attains a test MSE of $0.0019$ and an $R^{2}$ of $0.98$. Figure 8 displays representative snapshots of the simulation runs.
Figure 8: Snapshots of Prediction Performance for Simulation Study 2.
The high predictive accuracy of the simulation runs indicates that constrained lasso can efficiently find a functional basis from a select few agents and determine the agent weights to emulate the ground-truth data patterns based on limited observations. This is in fact not surprising because in the initial assignment of the endowments into the ground truth and proxy groups we have used a random partition. As informed by our proposed behavioral preference model, this causes the ground truth vectors—used in the simulation study to form the aggregate preference—and proxy vectors to span the same preference space. As the preference space is not directly accessible, we can derive a partial gauge through the observed agent responses. If the responses of a group of agents have higher variability, it is an indication that the preference space they span is large. Conversely, if agents tend to agree on their responses to different questions, it is a signal that their associated vectors in the preference space might overlap in sub-dimensions, leading to a pool generalization ability.
To effectively measure response variability, in the paper we introduce the notion of question entropy:
$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}$ (5)
where $H_{i}(\mathcal{A})$ is the normalized entropy of question $i$ based on responses from a group of agents $\mathcal{A}$, $K_{i}$ denotes the number of unique response options $\{1,\dots,K_{i}\}$ for question $i$, and $p_{ik}$ represents the empirical proportion of responses selecting option $k$. The normalization factor ensures comparison across questions. A high question entropy indicates that the responses are highly varied for this question.
For the simulation studies, we define group entropy as the average question entropy computed using responses from that group:
${H}(\mathcal{A})=\frac{1}{N}\sum_{i=1}^{N}H_{i}(\mathcal{A}).$ (6)
The group entropy is calculated using responses to all survey questions, regardless of the split. Based on it, we introduce the notion of Entropy Coverage Ratio (ECR) defined as the ratio between the group entropy of proxy agents and that of the ground-truth agents:
$ECR=\frac{{H}(\mathcal{A}_{\text{proxy}})}{{H}(\mathcal{A}_{\text{gt}})}.$ (7)
An ECR lower than 1 indicates that the proxy agents have less response variability than the ground truth, while an ECR greater than 1 indicates the opposite.
Figure 9: Entropy Coverage Ratio Diagnostics for Simulation Study 2.
The left panel of Figure 9 shows the distribution of entropy coverage ratios (ECRs) for the two analyzed scenarios. The mass is concentrated around $1.0$, corroborating the earlier observation that random partitioning tends to preserve entropy structure. The right panel plots ECR against the number of proxies for the proxy-subsampling scenario. The relationship between ECR and the number of proxies is subtly positive: when the number of proxies is small, the likelihood of drawing an unrepresentative sample increases. As the sample size grows, group-level entropy converges toward the population-level entropy—an empirical manifestation of the law of large numbers.
A.3 Response Variability Matters
In this final simulation study, we analyze the effect of the entropy coverage ratio on predictive accuracy for constrained lasso.
Figure 10 displays the group entropies for THE W42 endowments organized by modes—we call them mode entropies in the endowment generation logic. We subsume them further into three different tiers according to the group entropy value: in total, we have 71 endowments in the low entropy tier, 92 in the mid entropy tier, and 137 in the high entropy tier.
Figure 10: Endowment Count per Mode (Sorted by Entropy).
To cover a wide range of ECR ratios, we design the following simulation strategy:
  1. 1.
    Ground Truth Construction. For each simulation round, pick one of the three tiers as the ground-truth tier and sample 30 endowments from the tier to construct the ground-truth agents and assign ground-truth weights. The unselected endowments from the tier are joined with the endowments from the other two tiers to form the pool of proxy candidates.
  2. 2.
    Proxy Agents Construction. For each simulation run, we select 30 proxy agents to form the proxy pool. We begin by selecting agents from the lowest entropy modes (ensuring that the cumulative candidate endowments surpass 30), and in each consecutive run, include the next mode with higher entropy into the selectable modes. We sweep across the modes by increasing mode entropy.
  3. 3.
    Proxy Agents Selection and Response Emulation. In each simulation run, after the proxy pool has been constructed, we use constrained lasso to select proxy agents to emulate the observed aggregate response data.
Figure 11: Mean Squared Error and Coefficient of Determination ($R^{2}$) by Entropy Coverage Ratio (ECR) for the trainval set (left) and test test (right). Entropy Coverage Ratio is defined as the ratio of the average response entropies of the proxies and ground-truth agents. LOWESS curves are fit using 10 repetitions.
We repeat each simulation 10 times to form the final results. Figure 11 visualizes the train and test performance of the constrained lasso by the entropy coverage ratio. As the ECR increases, the predictive accuracy improves at both the training and the test times, lending further credence to the analysis based on the preference model in the second simulation study.
Figure 12: Snapshots of Prediction Performance for Simulation Study 3.
Figure 12 offers three snapshots of the simulation runs corresponding to ECRs equaling $0.50$, $1.01$ and $1.67$. When proxy entropy is low compared with ground truth, the proxy agents fail to provide a sufficiently expressive basis to span the aggregate ground-truth preference. As a result, constrained Lasso can only approximate a lower-dimensional projection of the true preference signal. This results in poor prediction performance, as illustrated in the leftmost snapshot. Conversely, when proxy entropy exceeds ground-truth entropy, prediction error begins to increase—albeit at an infinitesimal scale. The subtle deterioration of performance when ECR surpasses 1 could be due to various reasons. A plausible cause is from the data perspective: when the proxy pool becomes too diverse, it may introduce spurious variability that does not align well with the true underlying structure. Put differently, an observed response might amalgamate signals from different latent preference factors, some of which are irrelevant to the aggregate preference in focus. While greater diversity expands the representational capacity of the proxy pool, it may also dilute the signal and induce overfitting, thereby reducing generalization accuracy.
The key takeaway from the simulation study is that absent the knowledge of ground truth, it is crucial to ensure an adequate group entropy among the proxy agents to form a functional basis for ground-truth emulation. However, one should also refrain from solely relying on entropy as a measure of preference diversity lest spurious correlations beguile us under limited data regimes. While our alignment method is intentionally designed to function with minimal data—requiring only aggregate response data—in practice, if individual human responses are available, a practitioner may use the group entropy computed from these responses as an anchor for endowment generation.
Appendix B Taking P2P to the Wild: Supplementary Experiment Results
B.1 Wave 42
Tracking entropy during active endowment generation
Figure 13 shows the entropy trajectories by question during active endowment generation. Of the tracked questions, 53 experience a clear rising trend in entropy, while 3 see a noticeable fall in entropy at the end of the generation, the other questions have stable entropy values across the update steps. Figure 14 displays the question entropies at the end of active endowment generation.
Figure 13: Entropy trajectories by question during active endowment generation. Trends are classified based on slope and volatility of the entropy trajectory: “rising” if the slope exceeds 0.01 and standard deviation is above 0.02; “falling” if the slope is below –0.01 with sufficient volatility; otherwise labeled “stable”.
Figure 14: Question entropies at the end of active endowment generation.
Lasso Selection
Figure 15 plots the cross-validated mean squared error (MSE) for each value of $\alpha$ examined during hyperparameter tuning. Among the 30 candidate values, cross-validation identifies the optimal $\alpha$ as $1.17\times 10^{-3}$. Figure 16 shows the number of agents selected by constrained lasso by mode entropy, together with the number of agents selected from each template/strategy. Mixed modes contribute the largest share to the ensemble with 20 proxy agents, followed by question-patching modes with 14. Thematic modes come next, contributing 5 agents, alongside 5 from the survey mode. The remaining 1 proxy is selected from the core mode.
Figure 15: Cross-validated MSE by alpha ($\alpha$), the penalization parameter, for constrained lasso. The best alpha selected by CV is $1.17\times 10^{-3}$.
Figure 16: Number of agents selected by constrained lasso by mode entropy. Mixed modes contribute 20 proxies to the ensemble, followed by question patching modes contributing 14. The thematic modes come third, contributing 5 proxies. Five proxy agents are selected from the survey mode, while the rest 1 is from the core mode.
Elastic Net Results
In the main paper, we show prediction results using constrained lasso. Here, we provide the results for constrained elastic net (Figures 17 and 18). Overall, the prediction results are similar those using constrained lasso.
Figure 17: Cross-validated MSE by alpha, the overall penalization parameter, and L1 ratio, the penalization weight for the L1 term, for constrained elastic net. The lowest cross-validated MSE is achieved at $\alpha=2.59\times 10^{-3}$, with l1 ratio set at $0.60$.
Figure 18: Prediction performance of the agent ensemble selected by constrained elastic net. Results are similar to those using constrained lasso.
B.2 Panel Study Across ATP Waves
To obtain a panoramic view of P2P’s performance empirically, we conduct a preliminary panel study applying P2P to a collection of 14 waves from the American Trends Panel (ATP). Each wave is repeated three times to assess the consistency and robustness of our method under varying random seeds and endowment draws. This subsection provides an extended empirical perspective on the real-world applicability of our approach and sheds light on variability patterns that may not arise in controlled simulation environments. The ATP waves used in this panel study are: W26, W27, W29, W32, W34, W36, W41, W42, W45, W49, W50, W54, W82, and W92.
Figure 19 illustrates the test MSEs for the 42 experiments organized by the number of survey questions per wave. As shown, within-wave performance is generally stable, with most waves exhibiting low variance in test MSE across repeats. However, test MSE varies more significantly across waves. The performance distribution is summarized below:
  • 8/14 waves achieve test MSE consistently below 0.015.
  • 4/14 waves fluctuate around 0.015.
  • 2/14 waves have higher error ranges (0.020–0.025).
These preliminary results reveal no clear relationship between test MSE and the number of survey questions, suggesting that wave-specific factors, such as survey content, attribute coverage, or question formulation, may play a larger role in shaping performance.
Figure 19: Test performance across ATP waves. Each point represents a wave-specific run, colored by wave identity. The horizontal dashed lines mark two threshold levels: 0.015 and 0.020, used to classify performance regimes.
To illustrate the interaction between question-level entropy dynamics and predictive performance, we include three snapshots of the panel study in Figure 20, showcasing the differences in question entropy trajectories during active endowment generation and prediction performance.
Figure 20: Representative snapshots for three ATP waves with different test-time MSE. Top: the worst-performing wave. Middle: a moderately performing wave. Bottom: the best-performing wave. Left: entropy trajectories classified as rising (green), falling (red), or stable (gray). Center/right: Lasso regression fits for aggregate response prediction. While rising entropy often coincides with improved fit, some waves achieve strong performance with limited entropy change, suggesting additional contributing factors beyond diversity expansion.
Interestingly, the wave with the lowest test MSE (bottom row) exhibits relatively modest entropy change, while the poorest-performing wave (top row) shows mostly flat entropy evolutions. This suggests that while entropy dynamics can be indicative of diversity expansion, strong initial endowment coverage or other wave-specific factors (e.g., question content or noise) may also influence alignment outcomes. Entropy trends are a useful diagnostic but should be interpreted in the context of wave-specific factors.
The patterns observed in the panel study warrant deeper analysis, which is likely to inform refinements that make P2P more stable and effective across settings. One especially promising direction is to extend the entropy-based active endowment generation strategy to support multi-objective criteria, allowing the system to balance diversity, coverage, and alignment more systematically.
Another promising extension is to include a module that directly measures question semantics. Based on our preference reconstruction theory, survey questions define the response space $\mathcal{R}$. Question entropy serves as a key signal for underlying diversity when conditioning on a fixed question. However, an equally important source of variation lies in the diversity across questions themselves. This includes differences in topic coverage, difficulty, and framing—factors that shape the expressiveness and completeness of the response space. Ideally, the question set should span a semantically broad topic space so that variations in response entropy meaningfully reflect differences in latent preferences. A dedicated semantic module would enable content-aware data-splitting strategies and support more rigorous generalization analysis across question types. This is an important direction for future development of P2P.
Due to budget constraints, our empirical evaluation relies on American Trends Panel (ATP) waves, which are topical in nature and tend to focus on a narrow set of issues per wave. While this allows for controlled testing, it limits the diversity of question types and themes. A more expansive and promising application domain for P2P is large-scale general social surveys, such as the General Social Survey (GSS) or World Values Survey, which span a wide range of topics and include high-quality, theoretically grounded questions relevant across disciplines. Testing P2P in such settings would provide a more comprehensive benchmark for its alignment and coverage capabilities, and yield better insights into its generalizability across domains of social inquiry. Moreover, the agent ensembles generated from general social surveys could, in turn, be used to provide feedback on smaller topical surveys to improve survey design and assess content quality.
Appendix C Methodological and Experimental Details
C.1 Prompts
C.1.1 Attribute Learner (AttributeLearner)
To generate attributes from a single question:
You are an intelligent research assistant trained to analyze individual survey questions and infer which human attributes might influence how different people respond.
You are given a single survey question. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to shape responses to this question.
Focus on underlying factors that would cause meaningful variation in answers across different types of people. Avoid generic or overly broad attributes.
Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
**Example output format:** ["religious affiliation", "political ideology", "trust in government"]
To generate attributes from a set of questions:
You are an intelligent research assistant trained to analyze survey questions and infer which human attributes might influence how individuals respond.
You are given a set of training-only survey questions. Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to affect responses to these questions.
Carefully analyze the content and framing of the questions. Identify underlying factors that might shape how different people respond. Focus on attributes that are salient, discriminative, and potentially variable across respondents.
Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
**Example output format:** ["attribute1", "attribute2", "attribute3", "attribute4"]
C.1.2 Endowment Model (EndowmentModel)
To generate endowments from attributes:
You are an expert assistant trained to generate realistic, diverse, and demographically plausible personas for social science surveys.
Each persona should include:
‘eid‘: a short, lowercase, variable-safe identifier that encodes key traits (e.g., urban_liberal_30s_female). No punctuation or spaces.
‘endow_text‘: a natural language description of the persona (1-2 sentences), written as if describing a survey respondent.
Instructions:
Represent a wide range of age, gender, race, education, region, and political ideology.
Avoid repetition of phrasing or demographic combinations across personas.
Do not include explanations or formatting outside of the JSON array.
C.1.3 Survey Conductor (SurveyConductor)
To use an agent to answer a survey question:
You are completing a survey.
Your answer should reflect the person described in the profile above, using their preferneces, beliefs and experiences.
Respond with only the final answer string, not the code or label in brackets.
Do not include any reasoning, explanation, or commentary.
Do not preface your answer with phrases like ’I would choose’.
Just return the answer text exactly as it appears in the options.
C.2 Active Endowment Generation
Algorithm 1 Active Endowment Generation
1: Initialize attribute bank $\mathcal{B}$ with core, thematic, and theoretical templates
2: Derive survey-specific attributes from $\mathcal{Q}_{\mathrm{train}}$ via AttributeLearner; append to $\mathcal{B}$ Initial Sampling Stage
3: Sample initial endowments $\mathcal{E}_{0}$ via equal-mode sampling from $\mathcal{B}$
4: Instantiate initial agents $\mathcal{A}_{0}$ from $\mathcal{E}_{0}$ and elicit responses $\mathcal{R}_{0}$ on $\mathcal{Q}_{\mathrm{train}}$
5: Initialize ThemeVariabilityTracker (henceforce, Tracker) with $\mathcal{A}_{0}$, $\mathcal{R}_{0}$, and $\mathcal{Q}_{\mathrm{train}}$. Expansion Stage
6:while agent budget $N_{A}$ not yet reached do
7:  Compute variability scores and mode sampling probabilities via Tracker
8:  Allocate endowment budget across modes according to probabilities
9:  if question patching is enabled then
10:   Identify lowest-entropy questions $\mathcal{Q}_{\mathrm{low}}$ via Tracker
11:   for each $q\in\mathcal{Q}_{\mathrm{low}}$ do
12:    if $q$ is queried for the first time then
13:     Extract attributes for $q$ via AttributeLearner and append to $\mathcal{B}$
14:    else if $q$ has appeared more than $n$ times then
15:     Enable mixed-mode strategy using $q$ and top-performing mode
16:    end if
17:   end for
18:  end if
19:  Generate new endowments $\mathcal{E}_{\mathrm{new}}$ using allocated budget over sampled modes, patched question modes (if any) and mixed modes (if any).
20:  Instantiate agents $\mathcal{A}_{\mathrm{new}}$ and elicit responses $\mathcal{R}_{\mathrm{new}}$ on $\mathcal{Q}_{\mathrm{train}}$
21:  Update: $\mathcal{A}\leftarrow\mathcal{A}\cup\mathcal{A}_{\mathrm{new}}$, $\mathcal{R}_{\mathcal{A}}\leftarrow\mathcal{R}_{\mathcal{A}}\cup\mathcal{R}_{\mathrm{new}}$
22:  Update Tracker with $\mathcal{A}$, $\mathcal{R}_{\mathcal{A}}$, and $\mathcal{Q}_{\mathrm{train}}$
23:end while
24:Return: Final agent pool $\mathcal{A}$ and full response set $\mathcal{R}_{\mathcal{A}}$
Algorithm 2 Tracker Update Procedure
1: Input: Current agent pool $\mathcal{A}$, current agent responses $\mathcal{R}_{\mathcal{A}}$, and training questions $\mathcal{Q}_{\mathrm{train}}$
2:for each question $i\in\mathcal{Q}_{\mathrm{train}}$ do
3:  Compute question entropy $H_{i}(\mathcal{A})$
4:end for
5:for each mode do
6:  Compute variability score $V(\mathcal{A}_{\mathrm{mode}})$
7:end for
8: Compute softmax probabilities from variability scores
9: Output: Variability scores, question entropies, and sampling probabilities.
C.3 Example Endowments
An example of an endowment for a progressive female in her 20s living in an urban area
[…]
A 24-year-old urban female who identifies as bisexual and holds a master’s degree in social work. She exhibits a high fairness bias and is often concerned about equality. Despite her awareness, she sometimes struggles with extreme response bias, influenced by her strong convictions.
An example of an endowment for a nonbinary in their 30s valuing liberty living in a suburban area
[…]
They are a 32-year-old nonbinary software developer living in the suburbs. They highly value personal liberty and have a strong work ethic, often putting in extra hours for their projects. With a master’s degree in computer science, they frequently engage with scientific research, feeling that understanding complex topics fosters innovation. Their income is above average, and they embrace intelligence as a vital asset for personal growth.
An example of an endowment for a white male progressive in his 60s living in a small town
[…]
Pat is a 67-year-old White male from a small town. He identifies as a progressive and believes strongly in the importance of intelligence in community discussions and personal growth. Pat, who identifies as heterosexual, is skeptical of the increasing influence of big corporations on local politics, and he often seeks media that highlights grassroots movements and community activism. He balances his media consumption between traditional outlets and local storytelling platforms, knowing the value of good journalism.
Appendix D Extended Literature Review
D.1 Large Language Models to Emulate Human Preferences
Leveraging the ability of LLMs to emulate human or human-group responses with considerable fidelity, LLM alignment or personalization Tseng et al. [2024] has been largely focused on emulating human responses using various techniques. These include learning from feedback or reward models Bai et al. [2022], Christiano et al. [2017], under distribution shifts Leike et al. [2018], Krueger et al. [2021], finetuning Tan et al. [2024], retrieval augmented generation Prahlad et al. [2025], representation learning Ren et al. [2024], and using human-generated Horton [2023], Castricato et al. [2025], Argyle et al. [2023] prompts. However, when diverse responses need to be generated from many users or user groups instead of individuals, the data and time resources required for alignment grow linearly with the number of LLMs that need to be aligned. To reduce alignment costs, some works instead rely on AI-generated Schuller et al. [2024], Salminen et al. [2024], Ge et al. [2025], Simmons [2023] prompts, or alignment to multiple user groups Zhao et al. [2024], Mondal et al. [2025], even entire populations Cao et al. [2025]. However, even the best models struggle with this task Cao et al. [2025]. Furthermore, responses from LLMs have been found to have less variation than responses from real surveys Bisbee et al. [2024], and LLMs have a tendency to respond in the middle category Wang et al. [2024]. Consequently, it is imperative to address these limitations and devise data- and time-efficient approaches that enable LLMs to more accurately and reliably serve as proxies for survey populations. Our work contributes to this goal by introducing an active learning–inspired alignment framework that dynamically constructs a diverse agent basis under resource constraints and reconstructs population-level preferences via regression-based aggregation, without relying on ground-truth demographic profiles.
D.2 Data Selection and Active Learning
To increase preference coverage while operating under budget constraints, our active endowment generation pipeline draws inspiration from machine learning paradigms, particularly active learning.
Traditional data selection methods aim to improve the data efficiency of supervised deep learning models by identifying the most informative or important training examples for generalization Coleman et al. [2020], Paul et al. [2021]. Selection via Proxy (SVP) Coleman et al. [2020] uses a lightweight proxy model to estimate sample utility at a fraction of the cost, while GraNd and EL2N scores Paul et al. [2021] identify important training examples early in training.
Similarly, deep active learning applies query strategies to iteratively select the most informative samples from a large pool of data to be added to the training dataset for retraining or fine-tuning Astorga et al. [2024], Hübotter et al. [2025], Li et al. [2025]. Active learning strategies balance exploitation, focusing on uncertain or high-impact samples, with exploration, seeking diverse and novel samples to broaden coverage.
In our system, we apply this logic to guide both endowment generation and question patching. Specifically, low-entropy responses are treated as indicators of low coverage or model uncertainty. We refer to question patching as a boosting strategy—not in the ensemble learning sense, but to distinguish it from the main adaptive sampling process—since it targets underperforming questions by allocating additional sampling to enhance expressivity. Meanwhile, the exploitation–exploration mechanism used to sample across existing modes reflects the same active learning principle, but with a different emphasis: prioritizing high-utility regions while maintaining representational diversity. Together, these components form an active learning–inspired loop that incrementally expands coverage of the preference space while minimizing redundant agent generation.
Appendix E Mathematical Formalization of the Preference Reconstruction Theory
E.1 Space Definitions and Mapping Framework
We define four key spaces for formalizaing our alignment framework:
  • Attribute space $\mathcal{B}\subset\mathbb{R}^{d}$: each $b\in\mathcal{B}$ is a structured configuration of interpretable attributes drawn from the attribute bank (e.g., ideology, altruism, personality). Attributes serve as control handles that steer LLM behavior.
  • Endowment space $\mathcal{E}$: each $e\in\mathcal{E}$ is a natural language prompt or agent description instantiated from an attribute configuration $b\in\mathcal{B}$. Endowments serve as the actual inputs to the LLM. We treat $\mathcal{E}$ as a structured set, not a Euclidean space.
  • Preference space $\mathcal{P}\subset\mathbb{R}^{m}$: each $p\in\mathcal{P}$ represents a latent vector encoding a behavioral inclination or preference profile. This space is unobservable but assumed to mediate how attributes shape responses.
  • Response space $\mathcal{R}\subset\mathbb{R}^{n}$: each $r\in\mathcal{R}$ is an observable vector of responses to survey questions (e.g., binary or Likert-coded).
Based on the four spaces, we define three mappings:
  • Attribute-to-Endowment Mapping ($\phi:\mathcal{B}\to\mathcal{E}$): This is a stochastic mapping that transforms a symbolic attribute configuration into a textual prompt or persona instantiation readily parsable by an LLM. Since multiple distinct prompts can be generated from the same attributes, we model $\phi$ as sampling from a conditional distribution over prompts given attributes.
  • Endowment-to-Preference Mapping ($g:\mathcal{E}\to\mathcal{P}$): This captures the internal LLM process by which a given prompt (endowment) induces a latent preference state. This mapping is governed by the LLM’s internal representations, shaped jointly by its large-scale pretraining corpora, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF). Albeit unobservable, this preference vector mediates the model’s behavioral tendencies in downstream response generation.
  • Preference-to-Response Mapping ($h:\mathcal{P}\to\mathcal{R}$): This final mapping generates a survey response vector based on the latent preference state. It represents how an agent’s inclinations are expressed as answers to survey questions.
E.2 Preference Reconstruction via a Functional Basis
The overall agent behavior function is thus a composition of the three mappings:
$f=h\circ g\circ\phi:\mathcal{B}\rightarrow\mathcal{E}\rightarrow\mathcal{P}\rightarrow\mathcal{R}$
where $f(b)$ denotes the response behavior of an LLM agent endowed based on attributes $b$.
Given a set of sampled agents $\mathcal{A}=\{b_{1},\dots,b_{k}\}$, we obtain a response basis $\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}$. We then treat alignment as a projection task:
$\min_{w\in\mathbb{R}^{|\mathcal{A}|}}\left\|\sum_{b\in\mathcal{A}}w_{b}f(b)-r_{\text{pop}}\right\|^{2}$ (8)
where $r_{\text{pop}}$ is the target aggregate response vector, and the weights $w_{b}$ reflect the inferred contribution of each proxy agent to the population preference.
E.3 Alignment as a Two-Stage Problem
We frame the alignment process as a two-stage problem grounded in mappings across attribute, endowment, preference and responses spaces.
Stage 1: Constructing a functional basis based on response variability.
The goal of the first stage is to construct a diverse and expressive set of LLM agents whose latent preferences collectively span the space where the ground-truth population preference resides. Since latent preferences are not directly observable, we use response variability as an indirect gauge, leveraging the preference-to-response mapping ($h:\mathcal{P}\to\mathcal{R}$). Starting from a sampled set of attribute configurations $\mathcal{A}=\{b_{1},\dots,b_{k}\}\subset\mathcal{B}$, we apply the composition mapping $f:\mathcal{B}\to\mathcal{R}$ to obtain a set of simulated agent responses:
$\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}\subset\mathcal{R}.$
We interpret variability in $\hat{\mathcal{R}}$ as an indirect signal of the latent variation in $\hat{\mathcal{P}}$. The design objective is to ensure that the span of $\hat{\mathcal{R}}$ is sufficiently rich so that the latent preference vectors form a functional basis for reconstructing population-level behavior. To optimize this objective, we operate in the more interpretable and tractable attribute space $\mathcal{B}$ rather than the high-dimensional preference space $\mathcal{P}$. By actively exploring semantically meaningful regions of $\mathcal{B}$—analogous to domain-aware active learning strategies Hao et al. [2024] and latent-manifold transformations Jin et al. [2024]—we adaptively improve coverage of $\hat{\mathcal{P}}$.
Stage 2: Projecting onto population responses.
In the second stage, we assume access to a population-level response vector $r_{\text{pop}}\in\mathcal{R}$, such as aggregate survey results. The construction of the agent ensemble is done by solving the optimization problem defined by (8), where each weight $w_{b}$ reflects the inferred contribution of the agent defined by attribute set $b$ in reconstructing the aggregate signal.
Together, these two stages—constructing a coverage-maximizing basis and estimating population-aligned weights—enable efficient and demographically-agnostic preference reconstruction using steerable LLM agents.

### v2

@@
 Abstract
-Large language models (LLMs) have demonstrated promise in emulating human-like responses across a wide range of tasks.
-In this paper, we propose a novel alignment framework that treats LLMs as agent proxies for human survey respondents, affording a cost-effective and steerable solution to two pressing challenges in the social sciences: the rising cost of survey deployment and the growing demographic imbalance in survey response data.
-Drawing inspiration from the theory of revealed preference, we formulate alignment as a two-stage problem: constructing diverse agent personas called endowments that simulate plausible respondent profiles, and selecting a representative subset to approximate a ground-truth population based on observed data.
-To implement the paradigm, we introduce P2P, a system that steers LLM agents toward representative behavioral patterns using structured prompt engineering, entropy-based sampling, and regression-based selection.
-Unlike personalization-heavy approaches, our alignment approach is demographic-agnostic and relies only on aggregate survey results, offering better generalizability and parsimony.
-Beyond improving data efficiency in social science research, our framework offers a testbed for studying the operationalization of pluralistic alignment.
-We demonstrate the efficacy of our approach on real-world opinion survey datasets, showing that our aligned agent populations can reproduce aggregate response patterns with high fidelity and exhibit substantial response diversity, even without demographic conditioning.
-Keywords Preference modeling  $\cdot$ Value alignment  $\cdot$ Evaluation
+Large language models are increasingly used as proxies for human subjects in social science research, yet external validity requires that synthetic agents faithfully reflect the preferences of target human populations.
+We introduce preference reconstruction theory, a framework that formalizes preference alignment as a representation learning problem: constructing a functional basis of proxy agents and recovering population preferences through weighted aggregation.
+We implement this via Prompts to Proxies (P2P), a modular two-stage system.
+Stage 1 uses structured prompting with entropy-based adaptive sampling to construct a diverse agent pool spanning the latent preference space.
+Stage 2 employs L1-regularized regression to select a compact agent ensemble whose aggregate response distributions align with observed data from the target population.
+P2P requires no finetuning and no access to sensitive demographic data, incurring only API inference costs.
+We validate the approach on 14 waves of the American Trends Panel, achieving an average test MSE of 0.014 across diverse topics at approximately 0.8 USD per survey.
+We additionally test it on the World Values Survey, demonstrating its potential to generalize across locales.
+When stress-tested against an SFT-aligned baseline, P2P achieves competitive performance using less than 3% of the training data.
+Preference Modeling, Value Alignment, Evaluation
 1 Introduction
-Pollsters are facing mounting challenges in the contemporary landscape, as public willingness to participate in surveys declines across regions and over time.
-In the United States, telephone response rates have dropped from 36% in the late 1990s to under 6% today Kennedy (2019), while major federal surveys such as the Current Population Survey (CPS) have experienced a 20-point drop since 2015 U.S.
-Bureau of Labor Statistics (2025).
-Likewise, across Europe Jabkowski and Cichocki (2025), response rates for national surveys in 36 European countries have fallen 0.8% per year on average between 1999 and 2018.
-This erosion of participation undermines statistical power and representativeness, particularly because the decline is not uniform across demographic groups.
-For instance, recent election polling failures have been traced to differential nonresponse: Republicans and independents are increasingly less likely to answer than Democrats Clinton et al.
-(2022); Parshall (2024).
-In the social sciences, this is a classic case of sample selection bias, which, if left unchecked, compromises both predictive accuracy and external validity of the survey-based inference.
-At the same time, the exorbitant cost of conducting large-scale surveys, due to respondent recruitment, compensation, and extensive piloting, places constraints on experimentation and iteration Kennedy (2015).
-These challenges highlight a pressing need for scalable, flexible, and theoretically grounded alternatives for modeling human preferences.
-The epistemic challenge of capturing human preferences, long studied in the social sciences, now finds a structural parallel in large language model (LLM) alignment, where researchers aim to steer model behavior in line with human values and beliefs.
-Standard approaches, such as supervised fine-tuning (SFT) Tan et al.
-(2024) and reinforcement learning from human feedback (RLHF) Bai et al.
-(2022); Christiano et al.
-(2017), train models to reflect normative ideals of helpfulness, harmlessness, or truthfulness.
-While effective for building general-purpose assistants, these techniques are inherently monolithic: they produce a single dominant behavioral profile, making them ill-suited for modeling pluralistic human populations.
-Scaling them to reflect multiple value systems requires retraining or fine-tuning for each group, rendering the approach costly, brittle, and inflexible.
-Consequently, these pipelines struggle with tasks that demand population-level diversity, such as simulating public opinion or generating group-sensitive predictions.
-To address this rigidity, alignment researchers have increasingly turned to post-training conditioning techniques that guide model behavior at inference time.
-A prominent example is the use of structured prompts, variously termed endowments Horton (2023), personas Castricato et al.
-(2025), or Silicon Samples Argyle et al.
-(2023), that encode sociodemographic attributes or behavioral dispositions.
-Prompted this way, LLMs can emulate individual-level responses with considerable fidelity Bisbee et al.
-(2024), making them attractive for constructing synthetic populations to pilot surveys or test hypotheses.
-However, this approach remains limited in key respects.
-First, it depends heavily on human-curated prompts or ground-truth demographic profiles, which may be unavailable or incomplete.
-Second, it often produces synthetic responses that lack the natural diversity of real populations: LLMs exhibit central-category bias Wang et al.
-(2024) and generate insufficient inter-agent variation Bisbee et al.
-(2024).
-Thus, while endowments reduce the need for retraining, they increase reliance on demographic templates and fall short of capturing true representational diversity.
-These limitations reveal a common assumption in alignment: that behavioral fidelity hinges on demographic accuracy.
-We challenge this view and propose a new theoretical framework grounded in preference reconstruction.
-Rather than aligning each synthetic agent to a ground-truth demographic profile, we argue that effective emulation requires only the recovery of the aggregate preference structure of the target population.
-Drawing on revealed preference theory in economics, we treat each agent as a proxy decison-maker whose responses expose latent beliefs or utilities.
-Crucially, we do not seek to recover the true basis of individual human preferences.
-Instead, it suffices to identify a functional basis in the latent preference space, a diverse set of proxy agents that span the opinion manifold, and determine appropriate aggregation weights over them.
-This abstraction is reflected in human societies: demographically distinct populations may exhibit similar collective attitudes, even if their constituent beliefs diverge.
-Our approach thus enables a more data- and alignment-efficient route to survey emulation, moving beyond identity-conditioned generation toward flexible distributional modeling.
-This perspective also clarifies and concretizes the goals of pluralistic alignment Sorensen et al.
-(2024), which advocates representing a spectrum of human values rather than optimizing for a singular objective.
-While compelling in principle, the pluralistic alignment literature has remained largely agnostic about how to measure pluralism or evaluate pluralistic systems.
-Our framework addresses this gap by proposing concrete criteria for measuring variation and fidelity, along with a mechanism to steer agent distributions efficiently under resource constraints so that the aggregate distribution of answers across the agent ensemble matches that of a given target population.
-In doing so, we operationalize distributional pluralistic alignment through a lens grounded in economic theory and suited for practical survey applications.
+Large language models (LLMs) hold considerable promise as proxies for human subjects in social science research.
+From agent-based modeling of community dynamics (Park et al., 2023; Zhao et al., 2024a), to simulating survey responses (Argyle et al., 2023; Park et al., 2024) and prototyping policy interventions (Hou et al., 2025), synthetic agents afford scalability, reproducibility, and cost advantages unmatched by traditional human samples (Horton, 2023; Manning et al., 2024).
+However, realizing this promise hinges on solving a fundamental validity challenge: the preferences expressed by synthetic agents must faithfully reflect those of the target human population, at least when restricted to the topic domain.
+Without such alignment, conclusions drawn from synthetic experiments lack external validity (Shadish et al., 2002), limiting the scientific value of LLM-based social simulation.
+This alignment challenge departs from conventional AI safety framing centered on steering a model toward a single set of desirable values, but shares motivation with the emerging notion of pluralistic alignment (Chen et al., 2025; Feng et al., 2024; Sorensen et al., 2024b) [^1] See Appendix E for an extensive review on this topic..
+The goal here is distributional fidelity—ensuring that an ensemble of synthetic agents reproduces the heterogeneity of preference patterns observed in a target population.
+In reality, this challenge is compounded by an invisible resource barrier: while well-resourced industrial labs can deploy supervised fine-tuning to train personalized models on large annotated corpora, small academic institutions and independent researchers often face limited compute, limited data, and ethical and practical hurdles that preclude access to sensitive personal information.
+In this regard, existing approaches offer unsatisfying trade-offs: off-the-shelf LLMs trained via supervised fine-tuning (SFT) (Tan et al., 2024) and reinforcement learning from human feedback (RLHF) (Bai et al., 2022; Christiano et al., 2017) are easy to deploy but tend to collapse toward central tendencies in their training data, obscuring the preference heterogeneity (Feng et al., 2024; Kirk et al., 2024; Lee et al., 2024; Slocum et al., 2025); demographic conditioning methods require sensitive personal data and labor-intensive profile construction, yet still fail to guarantee distributional fidelity (Argyle et al., 2023; Castricato et al., 2025; Park et al., 2024; Bisbee et al., 2024).
+In this work, we address this gap by introducing preference reconstruction theory, a framework that formalizes population preference alignment as a representation learning problem, inspired by revealed preference theory in economics.
+The core insight is that population-level preferences can be reconstructed by appropriately weighting a functional basis of proxy agents.
+We focus specifically on social survey data as our alignment signal, as they are publicly available, transparently documented, and designed to represent well-defined populations, yet typically withholding sensitive demographic information that would enable individual-level profile matching [^2] We expect the theory to extend to proprietary data in the format of categorical responses..
+We operationalize this theory via P2P, an inference-based alignment system with a modular two-stage design.
+Stage 1 uses prompt engineering and entropy-based adaptive sampling to construct a diverse pool of agents with different response patterns, aimed to span the latent preference space of a topic domain.
+Stage 2 employs L1-regularized regression to select a representative and compact ensemble of agents whose response distributions match those of the target human population.
+This modularity allows each stage to be adapted or extended independently, accommodating different sampling strategies or aggregation methods as needed.
+We validate the design through simulation studies and evaluate P2P on the American Trends Panel and World Values Survey, demonstrating robust performance across topics and locales at low cost.
 Our contributions are as follows:
-  • Theoretical foundation.
-We formalize preference alignment as a two-stage process: constructing a functional basis of proxy agents and determining aggregation weights, drawing on revealed preference theory to recover population-level preferences.
+  • Problem formulation.
+We formalize population preference alignment as a problem of representation learning, and introduce preference reconstruction theory, which states that learnable population preference can be represented using a functional basis of proxy agents with appropriately determined aggregation weights.
   • Modular alignment system.
-We implement this theory via a modular system, Prompts to Proxies (P2P), that combines active endowment generation (to promote response diversity) with regression-based aggregation (to estimate preferences without demographic alignment).
-  • Pluralistic alignment in practice.
-We provide the first operational framework for pluralistic alignment, introducing metrics for diversity and fidelity and propose scalable strategies to balance representational coverage and resource efficiency.
-  • Bridge to empirical social science.
-Our framework enables social scientists to simulate and test preference theories by controlling agent endowments and tracking their effects on response distributions.
-It also supports survey design and can mitigate nonresponse bias, offering a cost-effective complement to traditional data collection.
-  • Testbed for alignment research.
-The modular design provides a controlled environment for evaluating prompting strategies, enabling alignment researchers to explore steerability and value-sensitive modeling at scale.
+We operationalize preference reconstruction theory via Prompts to Proxies (P2P), a two-stage system that combines entropy-based adaptive sampling for diverse agent generation with L1-regularized regression for compact ensemble selection, requiring no fine-tuning, no demographic data, and modest resources to run.
+  • Empirical validation.
+We evaluate P2P across 14 ATP waves and the World Values Survey, showing 43% improvement over prompting baselines, cross-locale generalization with theoretically grounded error analysis, and competitive fidelity against SFT approaches at less than 3% of the training data.
 2 Modeling Preferences via Functional Bases
-Preference modeling under conventional lenses
-Human preferences are latent, complex, and often inconsistent, posing a persistent challenge for formal modeling.
-Economists and social scientists have approached this challenge through multiple modeling frameworks.
-One prominent approach models humans as rational agents who maximize a stable utility function.
-While elegant in theory and mathematically tractable, these models often struggle to explain empirical anomalies in observed behavior Kahneman and Tversky (1979); Thaler (1999); Kahneman and Tversky (2000).
-A complementary framework, revealed preference theory Samuelson (1938); Mas-Colell et al.
-(1995), shifts the focus from explicit utility functions to observable choices.
-It asks whether a pattern of choices can be rationalized as if generated by a utility-maximizing agent without assuming a specific functional form.
-This idea has proven especially valuable in empirical settings, offering a way to assess whether behavior is consistent with stable preferences directly from observed data.
-We take inspiration from the revealed preference perspective—not in its full rationality axioms, but in its core idea: preferences, though hidden, can be inferred from behaviors, such as responses to survey questions.
-This principle also underlies modern alignment pipelines such as SFT and RLHF, where models are steered toward human-preferred outputs using behavioral signals in the form of labeled data or feedback on generated responses.
-Our framework
-To formalize our approach, we zoom out from individual preferences and return to the level of survey analysis.
-Surveys are not designed to recover each respondent’s utility function; rather, they aim to capture aggregate population preferences, often through summary statistics such as response distributions or marginal means.
-This shift in perspective from individual to population invites a functional view (Figure 1): we can represent the population preference as a linear combination of individual preference vectors in a latent preference space.
-Responses to survey questions can then be viewed as mappings from this latent space to an observable response space.
-Crucially, this mapping is generally many-to-one: different underlying preferences may yield identical observed responses.
-Yet, as the number and diversity of responses increase, the expressive capacity of the response space improves.
-In the limit, a sufficiently rich battery of responses enables an increasingly precise identification of latent preferences.
-This perspective motivates our framework: Instead of recovering the ground-truth preference of each individual, we can construct a functional basis, a diverse set of proxy agents whose preferences span the latent space.
-By observing their responses and estimating the appropriate aggregation weights, which can be positive or negative, we can reconstruct the aggregate preference structure of the population.
+Alignment as representation learning
+We formalize human-AI preference alignment as a representation learning problem, inspired by the revealed preference theory (Mas-Colell et al., 1995; Samuelson, 1938).
+Let $\mathcal{P}$ denote the latent human preference space.
+An individual human preference is represented as a vector in this space, $p\in\mathcal{P}$, which gets revealed via a decision-making mechanism $h$ into observed responses, denoted by vector $r\in\mathcal{R}\subset\mathbb{R}^{n}$.
+Similarly, let $\widehat{\mathcal{P}}$ be the latent model preference space.
+[^3] We intentionally define the model preference space $\widehat{\mathcal{P}}$ as an abstract space to accommodate the fact that both training and prompting can lead to different model preferences.
+In this paper, we use model and LLM interchangeably, as we contextualize alignment in the NLP context.
+A model preference is represented as $\hat{p}\in\widehat{\mathcal{P}}$, which manifests through the inference mechanism $f$ into the same observed response space $\mathcal{R}$.
+The representation learning process can thus be abstracted as:
+[…]
+where $f^{-1}$ is implemented via a learning algorithm [^4] Finding $\hat{p}$ such that $f(\hat{p})\approx r$, not a literal inverse.. We introduce learnability of preference in a revealed sense.
+Definition 2.1.
+Individual human preference is learnable in the revealed sense if $)h(\mathcal{P})\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)).
+The definition using convex hull accommodates cases where individual human preference is not learnable by a single model but can be learned using an ensemble.
+Learning population preference via survey data
+While the above describes the learning process of an individual preference, alignment of large language models (LLMs) is usually done using responses from a human population.
+However, this target population is typically ill-defined for general-purpose LLMs, which have been through distinct stages of pretraining, supervised finetuning (SFT) and reinforcement learning from human feedback (RLHF), with data sources from different user groups.
+In this paper, we focus on preference alignment using aggregate responses in general social surveys, because these surveys are readily available, transparently documented, and designed to represent the views of well-defined target population.
+Formally, aggregate responses in a survey are defined as weighted averages of individual responses:
+$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}h(p_{i})\ ,$ (2)
+where $w_{i}$ is the weight of individuals with preference $p_{i}$ in the population, $I$. Accordingly, we extend our learnability definition to cover human population preference.
+Definition 2.2.
+Human population preference is learnable in the revealed sense if $)\mathrm{conv}\left(h(\mathcal{P})\right)\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)).
+It follows readily that human population preference is learnable in the revealed sense if and only if individual human preference is learnable in the revealed sense (cf.
+Proposition A.4).
+Hence, in this paper we focus the discussion on population preference, with the expectation that the results extend to individual cases.
+[…]
+Observation 2.3.
+If human population preference can be learned by a representative LLM agent, it can also be learned using a non-trivial [^6] A representative agent is by definition a degenerate ensemble.
+ensemble of LLM agents.
+Observation 2.4.
+If human population preference is learnable in the revealed sense, there could be cases where it can be learned by an ensemble but not by demographic conditioning nor a representative agent.
+Observation 2.5.
+If human population preference can learned by demographic conditioning, the ensemble that can learn it is not unique.
+Preference reconstruction theory
+Inspired by Observations 2.3-2.5, we propose preference reconstruction theory: instead of trying to find a one-to-one map between $p_{i}$ and $\hat{p}_{i}$, we can construct a functional basis—a set of proxy LLM agents with distinct preferences $\{\hat{p}_{j}\}_{j=1}^{J}$ spanning the relevant dimensions of human preferences—and then recover the human population preference through appropriate selection and weighting of these agents based on observational data.
+Figure 1 provides a functional perspective on preference reconstruction theory, together with a geometric interpretation of when the learnability assumption is violated (inadequate basis).
+Alignment as a two-stage problem
+Preference reconstruction via a functional basis can be formulated as a concrete learning problem:
+$\min_{\{\hat{p}_{j},\tilde{w}_{j}\}_{j=1}^{J}}L(r_{\mathrm{pop}},\sum_{j=1}^{J}\tilde{w}_{j}f(x_{0};\hat{p}_{j}))\ ,$ (3)
+where $L$ is a loss function.
+Directly optimizing (3) is challenging, especially as the dimension $J$ is an optimizer per se.
+As a solution strategy, we treat alignment as a two-stage problem with distinct stage goals.
+The first stage prioritizes the construction of a diverse set of proxy agents, measured in a revealed sense by their response variability.
+This goal aligns with the notion of pluralistic alignment (Sorensen et al., 2024b), which stresses the importance of representing a spectrum of human values rather than converging on a single ideal.
+The second stage concentrates on selecting a parsimonious ensemble from the set of candidate agents, based on the fit between ensemble aggregate responses and human ground truth.
+To avoid overfitting and test whether the ensemble has indeed learned the population preference, we partition survey questions into training and test sets, using the training set for selection and test set for validation.
+This solution strategy allows us to implement preference reconstruction modularly.
 Figure 1: Top: Finding a functional basis to recover latent ground-truth preference. Bottom: Identification of latent preference through multiple observations.
-Alignment as a two-stage problem
-Based on our theoretical framework, we can cast alignment as a two-stage problem.
-The first stage involves constructing a broad and expressive candidate basis of proxy agents, each representing a distinct vector in the latent preference space.
-For the basis to be effective, it must capture meaningful variation across the population.
-This objective aligns closely with recent calls for pluralistic alignment Sorensen et al.
-(2024), which stresses the importance of representing a spectrum of human values rather than converging on a single ideal.
-However, we do not have direct access to individual preferences, and the prompt space is far too vast to explore in an unstructured manner.
-To address this, we adopt a structured prompting strategy based on attributes—interpretable factors whose variation is likely to influence preference expression.
-These attributes may be informed from theory (e.g., ideology, values, levels of altruism) or content (e.g., topical relevance).
-By combining and varying these attributes systematically, we generate a set of endowment-conditioned agents that span different positions in the preference space.
-The second stage concentrates on designing an aggregation mechanism to reconstruct the population preference from proxy agents.
-We frame this as a supervised learning problem: given aggregate responses for a subset of survey questions, we estimate a set of aggregate weights over the proxy agents based on their individual response patterns.
-These weights are then used to predict responses on held-out questions, enabling out-of-sample evaluation.
-This setup allows us to examine whether the weighted agentic ensemble indeed captures the target population preference, rather than overfitting to the training set.
-We adopt regression-based methods for this purpose, as they align naturally with our assumption that population preferences emerge as linear combinations of individual vectors in the latent space.
-We describe implementation details, including attribute construction, endowment generation, and constrained regression methods, in the following section.
 3 Prompts to Proxies (P2P)
-To implement our preference reconstruction strategy in the survey context, we engineer a modular alignment system called Prompts to Proxies (P2P) (Figure 2).
-The system consists of two core components: (1) active endowment generation, powered by a dynamic attribute bank, and (2) regression-based aggregation, where constrained variable section algorithms form the backbone.
-They concretize the two stages of alignment introduced in the previous section and enable the generation of a compact set of opinionated LLM agents to emulate the results of observed population-level surveys.
+To operationalize our preference reconstruction strategy in the survey context, we engineer a modular alignment system called Prompts to Proxies (P2P) (Figure 2).
+The system consists of two core components: (1) active endowment generation, powered by a dynamic attribute bank, and (2) regression-based aggregation, where constrained a variable selection algorithm forms the backbone.
+They concretize the two stages of alignment introduced in the previous section and enable the generation of a compact set of opinionated LLM agents to emulate population-level survey results.
+3.1 Survey Splitting and Evaluation Protocol
+To support the alignment procedure, the full set of survey questions denoted $\mathcal{Q}$ are partitioned into training, validation, and test subsets $\{\mathcal{Q}_{i}\}_{i=\mathrm{train},\mathrm{valid},\mathrm{test}}$.
+$\mathcal{Q}_{\mathrm{train}}$ is used in Stage 1 while $\mathcal{Q}_{\mathrm{train}}$ and $\mathcal{Q}_{\mathrm{valid}}$ are used in Stage 2.
+$\mathcal{Q}_{\mathrm{test}}$ serves as the validation to test preference generalizability.
 Figure 2: System Overview of the Prompts to Proxies (P2P) Framework.
@@
 An external validity check accesses how well the weighted ensemble predicts aggregate responses on held-out questions, assessing out-of-sample generalization.
-3.1 Survey Splitting and Evaluation Protocol
-To support the alignment procedure, we partition the full set of survey questions, denoted $\mathcal{Q}$, into three disjoint subsets: $\mathcal{Q}_{\mathrm{train}}$, $\mathcal{Q}_{\mathrm{valid}}$, and $\mathcal{Q}_{\mathrm{test}}$, corresponding to the training, validation, and test sets, respectively.
-Let $N=|\mathcal{Q}|$ denote the total number of questions, and $N_{\mathrm{train}}$, $N_{\mathrm{valid}}$, and $N_{\mathrm{test}}$ the cardinalities of the three subsets.
-P2P allows users to set the split sets and ratio flexibly.
-For the alignment system, $\mathcal{Q}_{\mathrm{train}}$ is used for active endowment generation.
-$\mathcal{Q}_{\mathrm{train}}$ and $\mathcal{Q}_{\mathrm{valid}}$ are used in regression-based aggregation.
-We reserve $\mathcal{Q}_{\mathrm{test}}$ as an ultimate test of the predictive accuracy of the selected agent ensemble, ensuring that the reconstructed preference has a decent degree of generalizability.
-For regression analysis, categorical questions are binarized into question-option pairs.
 3.2 Attributes as Control Handles
-As a crucial part of the active endowment generation logic, we use attributes to shape model preferences.
-This approach, though counter-intuitive at first glance, is in fact a deliberate reverse engineering design choice.
-In humans, attributes such as ideology, dispositions, or values are better understood as reflections or downstream manifestations of latent preferences.
-However, language models do not possess preferences in the human sense.
-As next-token prediction machines, their behavior is shaped by statistical associations learned from human text corpora.
-This distinction allows us to invert the causal direction: rather than treating attributes as reflections of internal preference states, we tap into the pattern matching nature of LLMs and use attributes as control handles to steer model behavior.
-To ensure that attributes effectively steer model preferences, we design a preset attribute bank that draws inspiration from theories across diverse social science disciplines, such as economics, political science, management science, and psychology.
-The bank is organized hierarchically: individual attributes are grouped into modes, each representing a coherent subject area (e.g., values and beliefs, cognitive biases, or political ideology).
+Language models, as next-token predictors, lack intrinsic preferences like humans.
+Their outputs are governed by statistical associations learned from text corpora and shaped by the user prompts.
+The prompt space is far too vast to explore in an unstructured manner.
+To address this, we adopt a structured prompting strategy based on attributes—interpretable factors, such as ideology, dispositions, or values, whose variation is likely to influence preference expression.
+P2P treats attributes as control handles for steering model behavior.
+[^7] This contrasts from the human case, where such attributes are typically viewed as reflections of latent preferences.
+P2P’s active endowment generation deliberately leverages this inversion: rather than inferring attributes from preferences, it uses attributes to shape the model’s expressed preferences.
+P2P constructs endowments (persona profiles) by drawing attributes from the attribute bank and instantiating each attribute with a specific value through a specialized LLM termed endowment model (cf.
+Appendix B.2.2) [^8] The use of endowment model ensures the created endowment remains coherent and interpretable to human readers..
+The preset attribute bank for P2P lists example attributes taken from different social science disciplines, such as economics and political science.
+The bank is organized hierarchically: individual attributes are grouped into modes, each representing a coherent subject area (e.g., cognitive biases, political ideology).
 These modes are further clustered into higher-level templates, which reflect different design purposes or theoretical orientations.
-We define three primary templates.
-The core template contains general demographic descriptors and serves as the default baseline.
-The thematic template includes topic-specific modes such as economics, politics, culture and history.
-The theoretical template groups attributes derived from canonical frameworks such as Maslow’s hierarchy of needs Maslow (1943) and the Big Five personality traits John and Srivastava (1999).
-A condensed summary of these templates, representative modes, and attributes is shown in Table 1.
-These templates and modes are not fixed: they serve as a structured starting point but are fully extensible.
-Users can define new templates, introduce domain-specific modes, or modify existing attribute sets to suit different experimental contexts.
-By sampling and varying attributes from this flexible bank, we create prompts that instantiate distinct proxy agents occupying different regions of the latent preference space.
-Template Representative Modes and Attributes
-Core core: gender, race, ethnicity, education, income level,
-Thematic economics: occupation, price sensitivity, wealth/debt,
-politics: policy preferences, civic engagement, ideological orientation,
-behavioral_biases: status quo bias, loss aversion, overconfidence bias,
-Theoretical maslow: safety needs, self-actualization, love and belonging needs,
-big_five: conscientiousness, neuroticism, extraversion,
-Table 1: Structured view of the attribute bank, with representative modes and attributes grouped by template. Users can extend the bank with extra templates, modes, or attributes.
-While the attribute bank provides structured coverage, we also enable dynamic attribute discovery.
-We introduce two additional freeform templates that are specific to the survey under analysis.
-The survey template derives attributes by prompting a designated LLM, called attribute learner, to identify relevant decision-making factors based on the overall survey topic covered by the training set of questions, $\mathcal{Q}_{\mathrm{train}}$.
-The question template goes further by asking the learner to derive attributes tailored to a specific survey question.
-In both cases, the inferred attributes are appended to the attribute bank only for the duration of the survey run.
-This hybrid strategy strikes a balance between theory-driven and data-driven endowment generations, and enables human-AI co-discovery of salient preference descriptors.
-With the attribute bank in place, we build an endowment generator using a dedicated LLM termed endowment model that takes a set of attributes as input and produces diverse agent profiles by varying the instantiation of each selected attribute.
-3.3 Active Endowment Generation
-To construct a diverse preference basis without exhaustively enumerating attribute combinations, we create an active endowment generation procedure to operate under resource constraints.
-It proceeds in iterative steps.
-At each step, new endowments are sampled, instantiated as agents to elicit responses to survey questions, and evaluated based on their contribution to response diversity.
-The remainder of this section describes the key components of this procedure, including our variability scoring strategy, adaptive sampling steps, and entropy-based patching mechanisms.
+Figure 4 offers a structured view of the attribute bank and the endowment generation workflow.
+While the attribute bank ensures structured coverage, we also enable dynamic discovery through two freeform templates that are specific to the survey under analysis.
+The survey template prompts an LLM attribute learner to extract decision-making factors from the overall survey topic.
+The question template derives attributes for a specific survey question.
+These attributes are added to the bank only for the survey run.
+This hybrid approach balances theory- and data-driven methods, enabling human-AI co-discovery of key preference descriptors.
+3.3 Stage 1: Active Endowment Generation
+Stage 1 of P2P constructs a diverse preference basis without exhaustively enumerating attribute combinations.
+It begins by invoking an attribute learner to infer survey-specific attributes from the training questions, augmenting the attribute bank.
+Endowment generation then proceeds via adaptive sampling by mode: After initializing with equal endowments per mode, subsequent steps follow a bandit-style strategy—prioritizing modes that yield high response variability while exploring new modes for under-represented question items (see full Algorithm 1 in Appendix B.3).
 Tracking response diversity with variability score
-As established in Section 2, observed responses can serve as behavioral manifestations for latent preference vectors.
-Through this link, response variability serves as an initial diagnostic for assessing diversity in the underlying preference space.
-To effectively measure response variability, we introduce the notion of question entropy:
-$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}$ (1)
+To measure preference diversity of agents in a revealed sense, we introduce the notion of question entropy:
+$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}\ ,$ (4)
 where $H_{i}(\mathcal{A})$ denotes the normalized entropy of question $i$ based on responses from a group of agents $\mathcal{A}$, $K_{i}$ is the number of unique response options $\{1,\dots,K_{i}\}$ for question $i$, and $p_{ik}$ is the empirical proportion of responses selecting option $k$.
 The denominator normalizes for scale, ensuring that entropy is comparable across questions with different numbers of choices.
-A high question entropy indicates greater response diversity for that question.
-By varying the group $\mathcal{A}$, e.g., restricting it to agents sampled from a particular mode, we can compute question entropy at the group level and compare the diversity contributes across modes.
-To turn question entropy into a useful metric for adaptive sampling, we introduce variability score for each mode. Formally, it is defined as
-$V(\mathcal{A}_{\mathrm{mode}})=\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\frac{H_{i}(\mathcal{A}_{\mathrm{mode}})}{H_{i}(\mathcal{A}_{\mathrm{total}})}.$ (2)
+A high question entropy indicates greater response diversity for that specific question.
+By restricting the group $\mathcal{A}$ to agents sampled from a particular mode (subject area), we can compute question entropy at the mode level, termed mode-question entropy, and compare the response variability across modes.
+Adaptive sampling references the variability score for each mode. Formally, it is defined as
+$V(\mathcal{A}_{\mathrm{mode}})=\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\frac{H_{i}(\mathcal{A}_{\mathrm{mode}})}{H_{i}(\mathcal{A}_{\mathrm{total}})}\ ,$ (5)
 where $H_{i}(\mathcal{A}_{\mathrm{mode}})$ denotes the entropy of responses to question $i$ based on the cumulative set of agents from a given mode up to the current generation step, and $H_{i}(\mathcal{A}_{\mathrm{total}})$ is the corresponding question entropy across all agents generated so far.
 Computed this way, the variability score favors modes that are capable of generating endowments which elicit diverse responses to questions that other modes typically fail to diversify.
-The intuition is to maximize the informational utility of each observation under budget constraints.
-When most agents agree on a question, a divergent response signals a valuable distinction in the underlying preference space.
-In P2P, we task a dedicated class object, Tracker, to compute the variability score for each mode after each generation round.
-The scores are then passed through a softmax function to produce a probabilty distribution over modes, which is used to guide endowment sampling in the next generation step.
-Adaptive sampling and update steps
-In P2P, endowment generation takes several steps.
-At the outset of the generation, an attribute learner is invoked to infer attributes from the survey based on the training questions, yielding the survey template on top of the existing templates in the attribute bank.
-The system then enters an initial sampling stage, drawing an equal number of endowments from each mode.
-The endowments are then used to instantiate agents to elicit responses to training questions, $\mathcal{Q}_{\mathrm{train}}$.
-The collective responses are then passed to Tracker for the computation of the variability scores and the sampling probabilities for the next update step.
-At the next update step, endowment budget allocation follows the multinomial distribution.
-A higher probability exploits the top-performing mode, while randomness allows exploration of its under-performing peers.
+In P2P, we task a dedicated class object, Tracker, to compute the variability score for each mode after each generation round (see Algorithm 2 in Appendix B.3).
+The scores are then passed to a softmax function to produce a probabilty distribution over modes, guiding endowment sampling in the next generation step via a multimonial draw—balancing exploitation of top-performing modes with exploration of under-performing ones.
 Patching low-entropy questions
-In addition to the existing trade-off between exploitation and exploration, we adopt an additional boosting strategy, which we call question patching.
+In addition to the existing trade-off between exploitation and exploration, we adopt an additional patching strategy, which we call question patching.
 At the end of each update step, as the Tracker computes the variability scores for each existing mode, it also keeps track of the questions with the lowest entropy values.
-When question patching is activated, the attribute learner will be called in the next generation step to exclusively infer attributes on each of the lowest-$n$ entropy questions, where $n$ is user-defined.
-The split of the endowment budget between the active generation and boosting generation is governed by a user-specified ratio.
-Figure 3: Entropy change across update steps for each question during active endowment generation for ATP W42.
-Each column represents a question.
-Green indicates an increase in entropy from the previous step, while red indicates a decline.
+When question patching is activated, the attribute learner will be called in the next generation step to exclusively infer attributes on each of the lowest-$k$ entropy questions, where $k$ is user-defined.
+The split of the endowment budget between active generation and patching generation is governed by a user-specified ratio.
 To avoid wasting resources on persistently low-entropy and to promote cross-pollination between modes, we additionally introduce a mixed mode strategy.
-It is triggered after a question appears $n$ times in the lowest-$k$ entropy list, where $n$ is a user-defined threshold.
-When in effect, the strategy relocates the assigned endowment budge to a new mixed mode, which combines attributes from the question with those of the current top-performing mode.
-This encourages diversity injection from high-performing modes while maintaining a question-specific focus.
+It is triggered after a question appears $t$ times in the lowest-$k$ entropy list, where $t$ is a user-defined threshold.
+When in effect, the strategy relocates the assigned endowment budge to a new mixed mode, which combines attributes from the question with those of the current top-performing mode, allowing diversity injection from high-performing modes while maintaining a question-specific focus.
+Figure 10 shows example endowments generated from a real run of P2P, illustrating how different templates and modes yield distinct profiles.
 The adaptive sampling cycle continues until the generation budget $N_{A}$ (an upper bound on the maximum number of agents allowed) is exhausted.
-By iteratively targeting low-diversity questions and expanding the attribute space, active endowment generation progressively increases response diversity while preserving coverage over previously explored regions of the preference space.
-3.4 Regression-Based Aggregation
-With the candidate proxy basis assembled, P2P enters the second stage: Regression-based Aggregation.
-The goal at this stage is to find a functional basis from the candidate vectors to reconstruct the observed preference.
-Depending on the endowment budget and the total number of trainable questions $\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}$, we could end up in a situation where there are more variables than observations.
+By iteratively targeting low-diversity questions, expanding the attribute space and promoting cross-pollination between modes, active endowment generation progressively increases response diversity while preserving coverage over previously explored regions of the preference space (cf.
+Figure 5 in Appendix B.4).
+3.4 Stage 2: Regression-Based Aggregation
+With the proxy basis assembled, P2P enters the second stage: regression-based aggregation.
+The goal at this stage is to find a functional basis from the proxy basis to reconstruct the population preference based on the observational data.
+Depending on the endowment budget $N_{A}$ and the total number of trainable questions $\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}$, we could end up in a situation where there are more variables (proxy agents) than observations (trainval questions).
 Consequently, a variable selection algorithm is vital.
-Besides, even with $N_{\mathrm{train}}+N_{\mathrm{valid}}>>N_{A}$, variable selection allows us to retain a parsimonious agent ensemble, saving inference costs on further simulations.
-To conduct variable selection, we consider regression methods with an added L1 penalization term.
-We can use either hold-out validation or cross-validation (CV) for hyperparameter tuning.
+Besides, even with $|\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}|>>N_{A}$, variable selection allows us to retain a parsimonious agent ensemble, saving inference costs on further simulations.
+We consider regression methods with an added L1 penalization term and use cross-validation (CV) for hyperparameter tuning.
 The objective of the learning problem can be stated as
@@
 When $\Omega$ subsumes an additional L2 norm, it becomes constrained elastic net.
-Our simulation studies indicates that constrained lasso faithfully recovers the ground-truth agents from a large candidate pool under an adequate number of observations (Appendix A.1), and it is potent in selecting a proxy ensemble to recover ground truth even when the number of proxy agents is small (Appendix A.2).
-Both results corroborate using it as a selection backbone.
-Our simulation study also investigates the relationship between entropy and prediction fit—we leave the details to Appendix A.3 for interested readers.
-4 Taking P2P to the Wild
-As a first empirical test of the alignment framework, we deploy P2P on real-world social survey data.
-Specifically, we use the American Trends Panel (ATP), a nationally representative U.S.
-survey conducted by Pew Research Center in discrete waves.
-In this demonstration, we choose Wave 42, which probes trust in science by asking respondents about scientific competence, integrity, and public accountability.
-The wave includes responses from approximately 4464 U.S.
-adults Pew Research Center (2019) and the cleaned dataset contains 128 questions.
-For the experiment, we split the questions according to a $7:1.5:1.5$ ratio into the training, validation and test sets.
-For the endowment generation stage, we set the endowment budget to be $300$, with $10$ endowments generated for each mode during initial sampling.
-Subsequently, the rest of the endowments are generated across 5 update steps.
-In each step, we allocate 75% of the endowment budget to question patching, where the lowest 3 questions are considered.
-We set the trigger threshold of the mixed mode strategy at 1, so that the budget is reallocated if a question enters the lowest-3 list the second time.
-At the end of the generation, 5 questions are patched, with additionally 5 mixed modes created.
-Endowment generation improves entropy
+To validate L1-regularized regression as a credible aggregation algorithm, we conduct extensive simulation studies for constrained lasso (Appendix C), where ground truth population preference is constructed using synthetic agents [^9] This allows us to create a clean environment where alignment result is not affected by the learnability condition (Def.
+2.2)..
+Our simulation studies yield three main findings.
+First, constrained lasso faithfully recovers the ground-truth agents from a large candidate pool under an adequate number of observations, including minority profiles with low weights, and only resorts to reconstruction when observations are extremely scarce (Appendix C.1).
+Second, when ground truth agents are removed from the candidate pool, constrained lasso remains effective in selecting a proxy ensemble to recover ground truth even when the number of proxy agents is small (Appendix C.2).
+Third, proxy ensembles with insufficient response entropy fail to recover ground-truth preferences, while those with excess entropy show only marginal deterioration, corroborating our choice of response variability as the optimization goal of Stage 1 of P2P (Appendix C.3).
+Together, these results provide constructive evidence for P2P’s design choices.
+4 Empirical Evaluation
+Having validated P2P’s design choices through simulation, we now evaluate it on real human survey data.
+We use the American Trends Panel (ATP) (Pew Research Center, ), a nationally representative U.S.
+survey conducted by Pew Research Center.
+We obtain the digitalized versions of these datasets from the OpinionsQA benchmark (Santurkar et al., 2023), which repurposed ATP responses to study value pluralism and misalignment in language models.
+We assess P2P across 14 ATP waves, spanning diverse socio-economic topics.
+We then evaluate locale generalizability on the World Values Survey, conduct error analysis and ablation studies, and stress-test against a distribution-calibrated model under topic shift.
+4.1 Experimental Setup
+For each survey, we split the questions according to a $7:1.5:1.5$ ratio into the training, validation, and test sets.
+To benchmark P2P, we create two prompting-based ensemble baselines with different endowment generation logic but feed both through the same regression-based aggregation stage as P2P for proxy selection and test set evaluation.
+This allows the baselines to benefit from the same data-driven selection mechanism, isolating the contribution of P2P’s endowment generation strategy.
+Vanilla baseline
+Endowments are generated using an LLM prompted [^10] The endowment generation prompts for P2P and the Vanilla baseline can be found in Appendix B.2.
+directly to generate diverse agent profiles conditioned only on the survey topic.
+PERSONA
+Endowments are sampled randomly [^11] We use random sampling as it preserves the representativeness of the dataset.
+In Appendix D.1.1, we provide baseline variants using K-means clustering as a pre-processing step.
+Clustering does not improve performance but increases cost.
+from the demographically conditioned PERSONA dataset (Castricato et al., 2025), which contains 1000 persona profiles derived from U.S.
+census data.
+For all methods, we set the endowment generation budget to $300$.
+For P2P, initial sampling generates $10$ endowments for each mode; active endowment generation is run in 5 update steps with the preset attribute bank, question patching (lowest 3; 75% endowment budget), and 10 attributes drawn for each endowment generation.
+Unless otherwise noted, all models use Gemini-2.0-flash as the backend [^12] See Appendix D.1.2 for raw model performance., and constrained lasso is used for Stage 2 selection.
+4.2 Main Results: Baseline Comparison on ATP Waves
+Table 1: Performance comparison across 14 ATP waves.
+Test MSE, entropy, and cost (USD) reported.
+P2P reports mean over 3 runs.
+Baselines report single runs, as repeated runs on W27, W32, W42, and W45 show negligible variance (MSE std << 0.005), confirming stability (Appendix D).
+$\Delta$ shows relative test MSE improvement over the best baseline.
+Vanilla PERSONA P2P
+Wave Topic # Qs MSE Ent. Cost MSE Ent. Cost* MSE Ent. Cost $\Delta$
+W26 Guns 78 .015 .50 \$0.5 .021 .37 \$1.0 $\mathbf{.010}$ .55 \$0.6 33%
+W27 Auto 96 .030 .40 \$0.6 .042 .40 \$1.3 $\mathbf{.022}$ .54 \$0.7 27%
+W29 Gender 77 .030 .45 \$0.5 .030 .44 \$1.0 $\mathbf{.016}$ .51 \$0.6 47%
+W32 Comm. 98 .020 .55 \$0.6 .026 .56 \$1.3 $\mathbf{.016}$ .62 \$0.7 20%
+W34 Biomed 67 .022 .43 \$0.5 .020 .39 \$0.9 $\mathbf{.011}$ .54 \$0.5 45%
+W36 Lead. 139 .033 .34 \$0.9 .035 .39 \$1.8 $\mathbf{.019}$ .48 \$1.0 42%
+W41 2050 90 .022 .45 \$0.6 .023 .48 \$1.2 $\mathbf{.013}$ .55 \$0.8 41%
+W42 Science 129 .034 .32 \$0.9 .038 .29 \$1.7 $\mathbf{.011}$ .43 \$1.0 68%
+W45 Misinfo 95 .016 .39 \$0.6 .028 .36 \$1.2 $\mathbf{.008}$ .44 \$0.7 50%
+W49 Privacy 98 .045 .41 \$0.6 .047 .41 \$1.3 $\mathbf{.021}$ .53 \$0.7 53%
+W50 Family 128 .027 .36 \$0.8 .024 .36 \$1.7 $\mathbf{.013}$ .44 \$0.9 46%
+W54 Econ 116 .022 .50 \$0.8 .026 .54 \$1.5 $\mathbf{.014}$ .57 \$0.9 36%
+W82 Global 104 .025 .52 \$0.7 .019 .48 \$1.4 $\mathbf{.008}$ .57 \$0.8 58%
+W92 Poli. 77 .025 .59 \$0.5 .031 .57 \$1.0 $\mathbf{.017}$ .61 \$0.6 32%
+Average .026 .45 \$0.7 .029 .43 \$1.3 $\mathbf{.014}$ .53 \$0.8 43%
+* PERSONA cost only covers the response elicitation stage—higher if persona generation tokens are included.
+As Table 1 shows, P2P consistently outperforms both baselines, achieving an average test MSE of 0.014 compared to 0.026 (Vanilla) and 0.029 (PERSONA)—a 43% improvement over the best baseline.
+For P2P, 12 out of 14 waves achieve mean test MSE (across 3 runs) below $0.02$, confirming generalization across topics without wave-specific tuning.
+The agent pool created using P2P also achieves higher diversity than the baselines, indicated by the higher average entropy (0.53 vs.
+0.43-0.45), corroborating our simulation finding that response diversity supports preference reconstruction (Appendix C.3).
+Notably, this diversity gain is cost-efficient, as P2P runs at approximately \$0.8 per survey, similar to Vanilla (\$0.7) and substantially cheaper than PERSONA (\$1.3, excluding persona generation overhead) [^13] The PERSONA pipeline (Castricato et al., 2025) requires resampling and re-inference, adding further API usage costs..
+Performance comparison is robust under other distributional metrics (1-JSD, EMD) (Appendix D).
+4.3 Locale Generalization: World Values Survey
+To probe whether P2P generalizes beyound U.S.
+populations and ATP-specific survey design, we evaluate it on World Values Survey (WVS) Wave 7 (Haerpfer et al., 2022), using cleaned data from Cao et al.
+(2025).
+We focus on three representative regions: the United States, Great Britain, and Hong Kong SAR.
+Under the same experimental protocol [^14] We pass the locale information to the endowment model to ensure regional relevance of the generated endowments., P2P achieves comparable performance on US and GB (MSE below $0.015$), with somewhat weaker results on HK (MSE around $0.024$) (Table 2).
+Table 2: P2P performance on WVS Wave 7 across three locales. Values are mean $\pm$ std over 3 runs. (Full details in Appendix D.)
+Metric US GB HK
+Entropy $\uparrow$ .572 $\pm$ .015 .571 $\pm$ .004 .548 $\pm$ .021
+Test MSE $\downarrow$ .015 $\pm$ .001 .014 $\pm$ .001 .024 $\pm$ .001
+Cost $\downarrow$ 1.76 $\pm$ .01 1.71 $\pm$ .00 1.74 $\pm$ .01
+Figure 3: Error Analysis of Hong Kong WVS Results.
+We compare UK-native and HK-native with HK-transfer (US endowments applied to HK via P2P Stage 2).
+(A) Ground truth divergence between populations ($r=0.848$).
+(B) Transfer outperforms native in overall performance, corroborating the functional basis perspective.
+(C) Error gap scales with cultural divergence, suggesting learnability limits (Definition 2.2).
+(D) Despite higher test MSE, HK-native outperforms transfer on culturally divergent questions, indicating that localization helps where cultures genuinely differ.
+4.4 Error Analysis: Transfer vs. Native Alignment
+One plausible explanation for the WVS performance gap is that LLMs trained primarily on English and Western-centric data better reflect Western populations (Durmus et al., 2024).
+Yet, we want to probe it further by connecting to our theoretical framework: How much of the gap is explained by insufficient agent diversity and can we characterize the learnability limits?
+To address these questions, we take US endowments, which have higher entropy, and apply them to HK via P2P Stage 2 by refitting regression weights.
+We call this HK-transfer and compare it against those in Table 2 (called US-native and HK-native).
+Figure 3 presents four complementary analyses.
+We define cultural divergence for each question-option pair as $|\mathrm{GT}_{\mathrm{US}}-\mathrm{GT}_{\mathrm{HK}}|$, the absolute difference in response proportions between populations.
+Panel (A) shows correlated but distinct US and HK ground-truth response patterns ($r=0.848$), with variation around the diagonal confirming cultural differences.
+Panel (B) shows that HK-transfer achieves lower test MSE than HK-native, corroborating the functional basis perspective—a more diverse agent pool facilitates preference reconstruction, even when generated for a different locale [^15] This result cuts against demographic conditioning approaches, which assume locale-matched profiles are necessary for faithful preference reconstruction..
+However, Panels (C) and (D) reveal important nuances.
+Panel (C) shows that prediction error scales with cultural divergence, offering an empirical handle on learnability limits (Definition 2.2).
+In culturally distinct dimensions, Gemini-based ensembles exhibit a clear bias toward US ground truth.
+As indicated by the trends, conditioning on locale information provides a mild mitigation.
+Panel (D) further corroborates this, showing that despite higher test MSE, HK-native outperforms HK-transfer on culturally divergent questions.
+The takeaway is that while cross-locale transfer can leverage diversity on average, locale-specific generation may capture culturally distinctive preferences that a foreign basis cannot fully reconstruct.
+The analyses also demonstrate P2P’s potential as a benchmark for model steerability and bias research (cf.
+Appendix F).
+4.5 Ablation Studies
+We conduct ablation studies on ATP Wave 42 to isolate component contributions. Full results are in Appendix D.3.
+Attributes
+Disabling question patching reduces entropy from 0.445 to 0.399 and increases test MSE from 0.010 to 0.011, while disabling both freeform templates (question & survey) further degrades MSE to 0.013.
+This suggests that human-AI co-discovery of preference descriptors yields meaningful performance gains.
+Varying the per-endowment attribute cap (10 vs.
+20) has minimal effect.
+Endowment budget
+Performance improves with budget (110-450 agents) but plateaus after 3 update steps (210 agents). A no-update baseline (110 agents) underperforms, validating the benefit of adaptive sampling.
+Model backend
+We test P2P’s performance on five model backends: GPT-4.1-mini, GPT-4.1-nano, Gemini-2.0-flash, Gemini-2.5-flash and a locally hosted Qwen.
+GPT-4.1-mini achieves the best entropy (0.471) and test MSE (0.009).
+Gemini-2.0-flash ranks second at one-quarter the cost.
+Smaller models (GPT-4.1-nano and Gemini-2.5-flash-lite) perform worse, consistent with our intuition that limited model capacity constrains the learnability (Definition 2.2).
+Regression-based aggregation
+Replacing regression with simple averaging of agent responses increases test MSE from 0.0104 to 0.0254, confirming that regression is essential for learning population preferences, not just ensemble compression.
+Switching loss weighting from option level to question level has minimal effect on W42 results.
+4.6 Stress Test: Comparison with an SFT-Aligned Model under Topic Shift
+P2P is designed as an inference-based alignment method for limited-data and limited-compute regimes, so our baselines focus on prompting-based methods under similar constraints.
+However, supervised finetuning (SFT) offers an alternative when more data is available.
+For instance, Cao et al.
+(2025) train a distribution-calibrated model on pooled WVS data from 46 countries (approx.
+8,400 training entries) to predict answer distributions conditioned on locale.
+We stress-test P2P on their protocol, which induces a topic shift: training on general attitudinal and religious-ethical items (Q1+Q2) and testing on political-cultural items (Q3).
+We apply P2P to three locales (the United States, Great Britain, and Hong Kong SAR), using only 172-185 local training entries and no finetuning.
+Despite using less than 3% of the training data and no SFT, P2P achieves comparable Jensen-Shannon Divergence (1-JSD) (0.71-0.75 vs.
+0.78), with matching or better majority-class prediction accuracy (MCPA) (0.47-0.64 vs.
+0.43).
+We interpret this as evidence that some preference structure transfers across topics though we do not assume such transfer is guaranteed.
+Studying cross-topic generalization remains an important direction for future work (Appendix F).
+Details on the stress test are presented in Appendix D.4.
+5 Conclusion
+This work formalizes preference alignment as a problem of preference reconstruction and operationalizes it as a two-stage process: first, constructing a functional basis of proxy agents; and then, recovering aggregation weights through supervised learning.
+We introduce P2P, a modular alignment system that implements preference reconstruction, combining entropy-based adaptive sampling with regression-based aggregation.
+Our simulation results corroborate P2P’s design choices.
+Empirical performance shows P2P can recover preference patterns observed in real human survey data while remaining competitive against SFT-aligned models despite using significantly less data.
+Future work can extend P2P to accommodate freeform textual output and aggregation, analyze its use as a benchmark for model steerability, devise re-anchoring strategies for non-stationary preference, and incorporate semantic and psychometric techniques for improved robustness and generalization (Appendix F).
+Our work offers a first step towards these endeavors and a practical platform for interdisciplinary research.
+Acknowledgments
+This work was partially supported by Google Cloud free trial credits and by discretionary research funds and compute resources from the GLOW.AI group, led by Prof.
+Bryan Low.
+The views and interpretations expressed in this paper are solely those of the authors.
+Impact Statement
+This work uses publicly available, anonymized survey datasets with sensitive identifiable information removed.
+While the large language model (LLM) backends have been trained to align with human values, they may still propagate social biases and stereotypes, and some generated endowment profiles may contain opinionated content.
+Our goal is methodological—studying preference reconstruction at the aggregate level rather than individual personalization—which reduces profiling risks.
+We see potential for positive impact on pluralistic alignment and principled use of AI in social science research, but stress that the agent ensemble is an auxiliary tool for understanding human choice patterns and should not replace humans in critical policy-making scenarios.
+Appendix A Modeling Preferences via Functional Bases: Proofs
+Proposition A.0 (Existence of a representative agent).
+Let $\widetilde{\mathcal{P}}$ be a latent preference space (regardless of human or model), with an individual preference denoted by $\tilde{p}\in\widetilde{\mathcal{P}}$.
+Let $g$ be the map from $\widetilde{\mathcal{P}}$ to the observed response space $\widetilde{\mathcal{R}}\subset\mathbb{R}^{n}$.
+Then the following two statements are equivalent:
+  1. 1.
+[…]
+  2. 2.
+    For any aggregate response $\tilde{r}_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}\tilde{r}_{i}$ with $\{w_{i}\}_{i=1}^{I}$ convex weights and $\tilde{r}_{i}=g(\tilde{p}_{i})$, there exists $\tilde{p}_{\mathrm{pop}}\in\widetilde{\mathcal{P}}$ such that $g(\tilde{p}_{\mathrm{pop}})=\tilde{r}_{\mathrm{pop}}$.
+Proof.
+First, we prove $(1)\implies(2)$.
+[…]
+[…]
+By the definition of image, it follows that there exists $\tilde{p}_{\mathrm{pop}}\in\widetilde{\mathcal{P}}$ such that
+$g(\tilde{p}_{\mathrm{pop}})=\sum_{i}^{I}w_{i}\tilde{r}_{i}\triangleq\tilde{r}_{\mathrm{pop}}.$
+Claim: $(2)\implies(1)$.
+[…]
+[…]
+[…]
+Remark A.1 (Sufficient condition for representative agent).
+If the decision-making mechanism $g$ maps a preference into a convex subset of a probability simplex $\Delta^{n-1}$, then any aggregate of individual responses is realizable by a representative agent.
+Proof.
+A convex subset of a probability simplex is by definition convex. The result follows from the theorem. ∎
+The implications of this lemma are two-fold:
+  1. 1.
+[…]
+  2. 2.
+[…]
+See 2.1
+Definition A.2.
+[…]
+See 2.2
+Definition A.3.
+[…]
+Proposition A.4.
+Human population preference is learnable in the revealed sense if and only if individual human preference is learnable in the revealed sense.
+Proof.
+The “only if” part follows directly from $h(\mathcal{P})\subset\mathrm{conv}(h(\mathcal{P}))$. Here we focus on the “if” part.
+By definition, a human population preference is defined as
+$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}h(p_{i}).$
+Since individual human preference is learnable in a revealed sense, i.e., $)h(\mathcal{P})\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)), for any $p_{i}$, there exists $\{w_{ik},\hat{p}_{ik}\}_{k=1}^{K_{i}}$ such that
+$h(p_{i})=\sum_{k=1}^{K_{i}}w_{ik}f(\hat{p}_{ik}).$
+Plug it back into the definition of human population preference and we get
+[…]
+The last step follows from re-indexing. Therefore, $)\mathrm{conv}(h(\mathcal{P}))\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)). ∎
+See 2.3
+Proof.
+By Proposition A.0, if human population preference can be learned by a representative LLM agent, then
+[…]
+For any $r_{\mathrm{pop}}\in\mathrm{conv}\left(h(\mathcal{P})\right)$, it follows from the definition of human population preference that
+$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}h(p_{i}):=\sum_{i=1}^{I}w_{i}r_{i}.$
+[…]
+It follows that
+$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}f(\hat{p}_{i})\in\mathrm{span}\left\{f(\hat{p}_{i})\right\}_{i=1}^{I}.$
+In words, we can construct an ensemble where each agent represents in preference a ground-truth individual in the human population (demographic conditioning).
+NB: Here we implicitly assume that the human population consists of more than one individual with distinct preferences.
+If the population consists of only 1 individual, then the representative agent and the ensemble coincide.
+∎
+See 2.4
+Proof.
+[…]
+$).\mathrm{conv}(h(\mathcal{P}))\subset\mathrm{conv}(f(\widehat{\mathcal{P}}$)). (9)
+[…]
+$r_{\mathrm{pop}}=\sum_{j=1}^{J}\tilde{w}_{j}f(\hat{p}_{j}).$
+In other words, human population preference can be learned using an ensemble. ∎
+See 2.5
+Proof.
+[…]
+$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}r_{i}=\sum_{i=1}^{I}w_{i}f(\hat{p}_{i}).$ (10)
+However, since $r_{1}\in h(\mathcal{P})\subset\mathrm{conv}(h(\mathcal{P}))$, we know there exists a finite set of $\{r^{*}_{k}\}_{k=1}^{K}$ with ${r}_{k}^{*}\in h(\mathcal{P})$ and convex weights $\{\lambda_{k}\}_{k=1}^{K}$ such that
+$r_{1}=\sum_{k=1}^{K}\lambda_{k}{r}_{k}^{*}.$
+Plug this into (10) and we get
+$r_{\mathrm{pop}}=\sum_{k=1}^{K}w_{1}\lambda_{k}{r}_{k}^{*}+\sum_{i=2}^{I}w_{i}r_{i}.$
+[…]
+In conclusion,
+$r_{\mathrm{pop}}\in\mathrm{span}\{f(\hat{p}_{1}^{*}),\dots,f(\hat{p}_{K}^{*}),f(\hat{p}_{2}),\dots,f(\hat{p}_{I})\}.$
+Appendix B Prompts to Proxies (P2P): Implementation Details
+B.1 Attribute Bank
+For the preset attribute bank (Figure 4), we define three primary templates.
+The core template contains general demographic descriptors and serves as the default baseline.
+The thematic template includes topic-specific modes such as economics, politics, culture and history.
+The theoretical template groups attributes derived from canonical frameworks such as Maslow’s hierarchy of needs (Maslow, 1943) and the Big Five personality traits (John and Srivastava, 1999).
+These templates and modes are not fixed: they serve as a structured starting point but are fully extensible.
+Users can define new templates, introduce domain-specific modes, or modify existing attribute sets to suit different experimental contexts.
+By sampling and varying attributes from this bank, we create prompts that instantiate proxy agents at different locations of the latent model preference space.
+Figure 4: Attribute Bank and Endowment Generation Workflow.
+The preset attribute bank is organized hierarchically by template (core, thematic, theoretical), mode (e.g., economics, politics, maslow, big five), and attribute (e.g., race, gender, ethnicity, urban/rural residency, risk aversion).
+Endowment is generated by a specialized LLM (EndowmentModel), which takes drawn attributes as input and outputs a profile with instantiation of these attributes.
+B.2 Model Prompts
+B.2.1 Attribute Learner (AttributeLearner)
+To generate attributes from a single question:
+Attribute Learner: Learning Attributes from A Single Question You are an intelligent research assistant trained to analyze individual survey questions and infer which human attributes might influence how different people respond.
+You are given a single survey question.
+Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to shape responses to this question.
+Focus on underlying factors that would cause meaningful variation in answers across different types of people.
+Avoid generic or overly broad attributes.
+Respond **only** with a Python-style list of double-quoted strings.
+Do not include any explanation, headers, or prose before or after the list.
+**Example output format:** ["religious affiliation", "political ideology", "trust in government"]
+To generate attributes from a set of questions:
+Attribute Learner: Learning Attributes from A Set of Questions You are an intelligent research assistant trained to analyze survey questions and infer which human attributes might influence how individuals respond.
+You are given a set of training-only survey questions.
+Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to affect responses to these questions.
+Carefully analyze the content and framing of the questions.
+Identify underlying factors that might shape how different people respond.
+Focus on attributes that are salient, discriminative, and potentially variable across respondents.
+Respond **only** with a Python-style list of double-quoted strings.
+Do not include any explanation, headers, or prose before or after the list.
+**Example output format:** ["attribute1", "attribute2", "attribute3", "attribute4"]
+B.2.2 Endowment Model (EndowmentModel)
+System prompt:
+Endowment Model: System Prompt You are an expert assistant trained to generate realistic, diverse, and demographically plausible personas for social science surveys.
+Each persona should include: - ‘eid‘: a short, lowercase, variable-safe identifier that encodes key traits (e.g., urban_liberal_30s_female).
+No punctuation or spaces.
+- ‘endow_text‘: a natural language description of the persona (1-2 sentences), written as if describing a survey respondent.
+Instructions: - Represent a wide range of age, gender, race, education, region, and political ideology.
+- Avoid repetition of phrasing or demographic combinations across personas.
+- Do not include explanations or formatting outside of the JSON array.
+To generate endowments based on attributes:
+Endowment Model: Endowment Generation Based on Attributes Generate {n} diverse persona(s) that vary meaningfully along the following attributes: {attr_string}.
+Each persona should reflect a distinct combination or value expression of these traits.
+Return a JSON array of dictionaries, each with: - ‘eid‘: a short, lowercase, variable-safe identifier - ‘endow_text‘: a brief natural-language description of the persona
+To generate endowments based on survey topics (also used for Vanilla baseline):
+Endowment Model: Endowment Generation Based on A Survey Topic Generate {n} diverse personas for a survey experiment.
+Each persona must include: - eid: short, lowercase identifier - endow_text: a short natural-language description The survey topic is: {topic} Ensure diversity across demographics.
+Return the result as a JSON array.
+B.2.3 Survey Conductor (SurveyConductor)
+To use an agent to answer a survey question:
+Survey Conductor: Answering A Survey Question You are completing a survey.
+Your answer should reflect the person described in the profile above, using their preferneces, beliefs and experiences.
+Respond with only the final answer string, not the code or label in brackets.
+Do not include any reasoning, explanation, or commentary.
+Do not preface your answer with phrases like ’I would choose’.
+Just return the answer text exactly as it appears in the options.
+B.3 Active Endowment Generation
+Algorithm 1 Active Endowment Generation
+1: Initialize attribute bank $\mathcal{B}$ with core, thematic, and theoretical templates
+2: Derive survey-specific attributes from $\mathcal{Q}_{\mathrm{train}}$ via AttributeLearner; append to $\mathcal{B}$ Initial Sampling Stage
+3: Sample initial endowments $\mathcal{E}_{0}$ via equal-mode sampling from $\mathcal{B}$
+4: Instantiate initial agents $\mathcal{A}_{0}$ from $\mathcal{E}_{0}$ and elicit responses $\mathcal{R}_{0}$ on $\mathcal{Q}_{\mathrm{train}}$
+5: Initialize ThemeVariabilityTracker (henceforce, Tracker) with $\mathcal{A}_{0}$, $\mathcal{R}_{0}$, and $\mathcal{Q}_{\mathrm{train}}$. Expansion Stage
+6:while agent budget $N_{A}$ not yet reached do
+7:  Compute variability scores and mode sampling probabilities via Tracker
+8:  Allocate endowment budget across modes according to probabilities
+9:  if question patching is enabled then
+10:   Identify lowest-entropy questions $\mathcal{Q}_{\mathrm{low}}$ via Tracker
+11:   for each $q\in\mathcal{Q}_{\mathrm{low}}$ do
+12:    if $q$ is queried for the first time then
+13:     Extract attributes for $q$ via AttributeLearner and append to $\mathcal{B}$
+14:    else if $q$ has appeared more than $n$ times then
+15:     Enable mixed-mode strategy using $q$ and top-performing mode
+16:    end if
+17:   end for
+18:  end if
+19:  Generate new endowments $\mathcal{E}_{\mathrm{new}}$ using allocated budget over sampled modes, patched question modes (if any) and mixed modes (if any).
+20:  Instantiate agents $\mathcal{A}_{\mathrm{new}}$ and elicit responses $\mathcal{R}_{\mathrm{new}}$ on $\mathcal{Q}_{\mathrm{train}}$
+21:  Update: $\mathcal{A}\leftarrow\mathcal{A}\cup\mathcal{A}_{\mathrm{new}}$, $\mathcal{R}_{\mathcal{A}}\leftarrow\mathcal{R}_{\mathcal{A}}\cup\mathcal{R}_{\mathrm{new}}$
+22:  Update Tracker with $\mathcal{A}$, $\mathcal{R}_{\mathcal{A}}$, and $\mathcal{Q}_{\mathrm{train}}$
+23:end while
+24:Return: Final agent pool $\mathcal{A}$ and full response set $\mathcal{R}_{\mathcal{A}}$
+Algorithm 2 Tracker Update Procedure
+1: Input: Current agent pool $\mathcal{A}$, current agent responses $\mathcal{R}_{\mathcal{A}}$, and training questions $\mathcal{Q}_{\mathrm{train}}$
+2:for each question $i\in\mathcal{Q}_{\mathrm{train}}$ do
+3:  Compute question entropy $H_{i}(\mathcal{A})$
+4:end for
+5:for each mode do
+6:  Compute variability score $V(\mathcal{A}_{\mathrm{mode}})$
+7:end for
+8: Compute softmax probabilities from variability scores
+9: Output: Variability scores, question entropies, and sampling probabilities.
+B.4 Demonstration of P2P: An Example Run on Wave 42
+To illustrate how P2P operates in practice, we present a detailed walkthrough of a single run on ATP Wave 42 (Trust in Science, 129 questions), tracing the system from the active endowment generation through final ensemble selection.
+B.4.1 Experimental Setup
+Questions are split according to a $7:1.5:1.5$ ratio into the training, validation, and test sets.
+We set the endowment budget to be $300$, with $10$ endowments generated for each mode during initial sampling.
+Active endowment generation is run in $10$ update steps with the preset attribute bank, question patching (lowest 3; 75% endowment budget), and 10 attributes drawn for each endowment generation.
+Gemini-2.0-flash is used as the model backend.
+B.4.2 Active Endowment Generation
 Figure 5 displays the entropy change for each question across the update steps.
@@
 Marginal gains in entropy are observed for the majority of questions, whereas some experience mild entropy drops.
-Appendix B provides question entropy trajectories in levels.
-The three questions that exhibit the largest drops in entropy are among the highest-entropy questions at the initial sampling stage.
-The evolution elucidates the tradeoff made by the active endowment generator encoded in the variability scores.
+Figure 6 displays the question entropies at the end of generation.
+Figure 5: Left: Entropy change across update steps for each question during active endowment generation for ATP W42.
+Each column represents a question.
+Green indicates an increase in entropy from the previous step, while red indicates a decline.
+Right: Entropy trajectories by question (line chart).
+Active endowment generation progressively increases response diversity.
+Trends are classified based on slope and volatility of the entropy trajectory: “rising” if the slope exceeds 0.1 and standard deviation is above 0.02; “falling” if the slope is below –0.1 with sufficient volatility; otherwise labeled “stable”.
+Figure 6: Question entropies at the end of active endowment generation.
+B.4.3 Regression-based Aggregation
+For regression analysis, categorical questions are binarized into question-option pairs.
 For the selection of the agent ensemble, we fit both a constrained lasso and a constrained elastic net, using CV to select the penalization parameters.
-Consequently, lasso selects 53 out of the 300 synthetic agents to form the agent ensemble, while elastic net selects 43.
-The training and testing performances are similar between the two methods.
+Consequently, lasso selects 58 out of the 300 synthetic agents to form the agent ensemble, while elastic net selects 107.
+Figure 7 shows the hyperparameter selected for lasso using cross-validation.
+Figure 7: Cross-validated MSE by alpha ($\alpha$), the penalization parameter, for constrained lasso. The best alpha selected by CV is $1.62\times 10^{-3}$.
 Figure 8: Prompts to Proxies: Emulation Results for ATP W42.
-Top: Train (left) and test (right) prediction performances for the binarized responses.
-Bottom: Human and agent (pre- and post- alignment) aggregated responses for the sample test question RQ4_F1Ae_W42.
-Prediction indicates feasibility
+Left: Train (left) and test (right) prediction performances for the binarized responses using constrained lasso.
+Right: Human and agent (pre- and post- alignment) aggregated responses for the sample test question RQ4_F1Ae_W42.
 Figure 8 shows the prediction performances of the agent ensemble selected with constrained lasso.
-The train-time root mean squared error (RMSE) is 0.08, with test RMSE slightly higher at 0.09, indicating adequate generalization ability of the functional basis.
-The bottom panel of Figure 4 presents a snapshot for the aggregated responses for a test set question, comparing the human ground truth with agentic emulations before and after regression-based aggregation.
+The train-time root mean squared error (RMSE) is 0.06, with test RMSE slightly higher at 0.08, indicating adequate generalization ability of the functional basis.
+The right panel of Figure 8 presents a snapshot for the aggregated responses to a test set question, comparing the human ground truth with agentic emulations before and after regression-based aggregation.
 Overall, the results are encouraging, especially considering that P2P constructs the endowments based solely on preset and learned attributes and is agnostic on the ground-truth demographic data.
-The panoramic view and future directions
-In Appendix B, we provide a preliminary panel study, applying P2P to 14 ATP waves, each repeated 3 times.
-Performance variation is relatively low within each wave but considerably higher across the panel.
-In 8 out of 14 waves, P2P consistently achieves a test MSE below 0.015 across all repeats.
-Four waves exhibit test MSEs fluctuating around 0.015, while the remaining two fall in the range of 0.020 to 0.025.
-While our simulation studies suggest that entropy coverage, the number of observations and the size of the proxy pool all influence emulation performance, the preliminary panel results signal intriguing directions for future study.
-Understanding the sources of variation in empirical performance—particularly in relation to both the similarities and the distinctions from simulation results—offers a promising lens on the dynamics of human–AI alignment.
-These insights can also inform refinements to P2P ’s modular architecture and motivate extensions of its functionality.
-P2P as a means to an important end
-At a broader level, our framework operationalizes pluralistic alignment Sorensen et al.
-(2024).
-Its structured use of attribute-based control enables the representation of diverse value systems, while its regression-based weighting mechanism supports distributional pluralism without enforcing demographic fidelity.
-Crucially, P2P demonstrates that steerability and pluralism are not separate challenges but deeply intertwined: the ability to steer model behavior through promptable attributes is what enables meaningful diversity in aggregate preference reconstruction.
-Although developed in the context of survey emulation, P2P generalizes naturally beyond surveys.
-The underlying preference reconstruction logic can be applied to other forms of observational data by discretizing continuous variables and binarizing categorical ones.
-Moreover, while this paper focuses on modeling aggregate population preferences, the framework could be adapted for individual-level alignment—for example, by learning proxy ensembles that emulate preferences of a single person across varied contexts.
-Beyond its methodological contributions, P2P affords a modular platform for empirical alignment.
-Our work is among the first to streamline the use of large-scale social survey data—not as textual corpora for language model pretraining, but as a structured, theory-informed substrate for steering models toward specific populations, cultural milieus, and value contexts.
-It supports systematic evaluation of prompt engineering techniques, attribute design strategies and sampling heuristics.
-Researchers can test how different endowment structures affect response diversity, and how entropy metrics reflect or obscure latent preference variation.
-These tools offer a bridge between high-level alignment goals and grounded behavioral evaluation.
-Finally, P2P offers a testbed for social scientists to probe and operationalize preference theories.
-By investigating how assumptions about cognitive style, value orientation, or belief structure affect response distributions and mode-level entropy, researchers can explore the explanatory power to theoretical constructs in a controlled but flexible setting.
-In this sense, the system fulfills the spirit of its name: P2P connects peer disciplines, translates prospects into products, carries ideas from pseudocode to paper, and moves pluralistic alignment toward practical needs.
-5 Conclusion
-This work establishes a theoretical foundation for viewing preference alignment as a problem of preference reconstruction.
-Inspired by revealed preference theory, our approach formalizes alignment as a two-stage process: first, constructing a functional basis of proxy agents using attribute-guided prompting; and then, recovering aggregation weights through supervised learning.
-In doing so, our work offers the first tractable implementation of pluralistic alignment, along with quantitative metrics to assess both diversity and fidelity.
-Beyond its theoretical contributions, this framework offers a practical platform for interdisciplinary research.
-In the social sciences, it can support survey design, question testing, and nonresponse mitigation, serving as a cost-effective complement to traditional data collection.
-For researchers in LLM alignment, the modularity of our system provides a controlled setting to evaluate prompt engineering strategies and their influence on agent diversity and downstream performance.
-More broadly, it lays the groundwork for agentic systems that enable value-sensitive modeling across disciplines.
-Notably, P2P repositions social survey data not as training material, but as a guiding signal for culturally grounded model alignment.
-Appendices
-Appendix A Simulation Studies
+Figure 9 provides a breakdown of the agent ensemble selected by lasso.
+Notably, preset, freeform, and mixed templates all contribute non-trivially to the final ensemble.
+Figure 10 provides three examples of endowments generated and selected by P2P for the run on ATP W42, tracing back to their attributes, modes, and templates.
+Additionally, results of constrained elastic net are shown in Figures 11.
+Overall, the prediction results are similar to those using constrained lasso.
+Figure 9: Breakdown of Selected Agents (58 in total) for W42 by Template. Preset, freeform, and mixed templates all contribute non-trivially to the final ensemble.
+Figure 10: Examples of endowments generated by P2P for a run on ATP W42.
+Figure 11: Left: Cross-validated MSE by alpha, the overall penalization parameter, and L1 ratio, the penalization weight for the L1 term, for constrained elastic net.
+The lowest cross-validated MSE is achieved at $\alpha=6.95\times 10^{-3}$, with l1 ratio set at $0.10$.
+Right: Prediction performance of the agent ensemble (107 selected agents) using constrained elastic net.
+Results are similar to those using constrained lasso.
+Appendix C Design Validation: Simulation Studies
 In order to gauge the validity of the proposed preference model and the factors affecting alignment performance, we conduct three simulation studies where we have the privilege of ground-truth knowledge.
@@
 Construction of the Ground Truth Data
-To construct the ground truth data, we sample a fraction of the endowments as the ground truth and randomly assign weights to these endowments, which represent the proportions of the associated personae in the ground truth population.
+To construct the ground truth data, we sample a fraction of the endowments as the ground truth and randomly assign weights to these endowments, which represent the proportions of the associated personas in the ground truth population.
 The ground truth data—percentages of individuals choosing difference options for a given question—are computed using weighted averages of responses provided by the ground-truth agents.
@@
 Therefore, for the sake of building intuitions, we focus on constrained lasso in the simulations and expect the results to be largely indicative of constrained elastic net’s performance.
-A.1 Recovering Ground Truth Agents with Constrained Lasso
+C.1 Recovering Ground Truth Agents with Constrained Lasso
 As a first test of constrained lasso as a variable selection algorithm, we aim to evaluate its ability to recover ground-truth agents from a broader candidate pool.
@@
 Lasso’s ability to recover the ground truth agents from the agent pool degrades with lower observation-to-agent ratio, but its predictive accuracy only drops mildly.
-A.2 Emulating Ground Truth Responses with Proxy Agents
+C.2 Emulating Ground Truth Responses with Proxy Agents
 While the first simulation exercise has readily allured to constrained lasso’s ability to find a functional basis to explain the observed data, in this second simulation study we extend the investigation by removing the ground-truth agents from the selectable agent pool.
@@
 As the sample size grows, group-level entropy converges toward the population-level entropy—an empirical manifestation of the law of large numbers.
-A.3 Response Variability Matters
+C.3 Response Variability Matters
 In this final simulation study, we analyze the effect of the entropy coverage ratio on predictive accuracy for constrained lasso.
@@
 While greater diversity expands the representational capacity of the proxy pool, it may also dilute the signal and induce overfitting, thereby reducing generalization accuracy.
+In practice, we posit that over-diversity is less likely than under-diversity, as existing studies using demographic conditioning (Argyle et al., 2023; Park et al., 2024) indicate LLM agents do not match closely with human ground truth.
+This justifies optimizing response variability in Stage 1 of P2P.
 The key takeaway from the simulation study is that absent the knowledge of ground truth, it is crucial to ensure an adequate group entropy among the proxy agents to form a functional basis for ground-truth emulation.
@@
 While our alignment method is intentionally designed to function with minimal data—requiring only aggregate response data—in practice, if individual human responses are available, a practitioner may use the group entropy computed from these responses as an anchor for endowment generation.
-Appendix B Taking P2P to the Wild: Supplementary Experiment Results
-B.1 Wave 42
-Tracking entropy during active endowment generation
-Figure 13 shows the entropy trajectories by question during active endowment generation.
-Of the tracked questions, 53 experience a clear rising trend in entropy, while 3 see a noticeable fall in entropy at the end of the generation, the other questions have stable entropy values across the update steps.
-Figure 14 displays the question entropies at the end of active endowment generation.
-Figure 13: Entropy trajectories by question during active endowment generation.
-Trends are classified based on slope and volatility of the entropy trajectory: “rising” if the slope exceeds 0.01 and standard deviation is above 0.02; “falling” if the slope is below –0.01 with sufficient volatility; otherwise labeled “stable”.
-Figure 14: Question entropies at the end of active endowment generation.
-Lasso Selection
-Figure 15 plots the cross-validated mean squared error (MSE) for each value of $\alpha$ examined during hyperparameter tuning.
-Among the 30 candidate values, cross-validation identifies the optimal $\alpha$ as $1.17\times 10^{-3}$.
-Figure 16 shows the number of agents selected by constrained lasso by mode entropy, together with the number of agents selected from each template/strategy.
-Mixed modes contribute the largest share to the ensemble with 20 proxy agents, followed by question-patching modes with 14.
-Thematic modes come next, contributing 5 agents, alongside 5 from the survey mode.
-The remaining 1 proxy is selected from the core mode.
-Figure 15: Cross-validated MSE by alpha ($\alpha$), the penalization parameter, for constrained lasso. The best alpha selected by CV is $1.17\times 10^{-3}$.
-Figure 16: Number of agents selected by constrained lasso by mode entropy.
-Mixed modes contribute 20 proxies to the ensemble, followed by question patching modes contributing 14.
-The thematic modes come third, contributing 5 proxies.
-Five proxy agents are selected from the survey mode, while the rest 1 is from the core mode.
-Elastic Net Results
-In the main paper, we show prediction results using constrained lasso.
-Here, we provide the results for constrained elastic net (Figures 17 and 18).
-Overall, the prediction results are similar those using constrained lasso.
-Figure 17: Cross-validated MSE by alpha, the overall penalization parameter, and L1 ratio, the penalization weight for the L1 term, for constrained elastic net.
-The lowest cross-validated MSE is achieved at $\alpha=2.59\times 10^{-3}$, with l1 ratio set at $0.60$.
-Figure 18: Prediction performance of the agent ensemble selected by constrained elastic net. Results are similar to those using constrained lasso.
-B.2 Panel Study Across ATP Waves
-To obtain a panoramic view of P2P’s performance empirically, we conduct a preliminary panel study applying P2P to a collection of 14 waves from the American Trends Panel (ATP).
-Each wave is repeated three times to assess the consistency and robustness of our method under varying random seeds and endowment draws.
-This subsection provides an extended empirical perspective on the real-world applicability of our approach and sheds light on variability patterns that may not arise in controlled simulation environments.
-The ATP waves used in this panel study are: W26, W27, W29, W32, W34, W36, W41, W42, W45, W49, W50, W54, W82, and W92.
-Figure 19 illustrates the test MSEs for the 42 experiments organized by the number of survey questions per wave.
-As shown, within-wave performance is generally stable, with most waves exhibiting low variance in test MSE across repeats.
-However, test MSE varies more significantly across waves.
-The performance distribution is summarized below:
-  • 8/14 waves achieve test MSE consistently below 0.015.
-  • 4/14 waves fluctuate around 0.015.
-  • 2/14 waves have higher error ranges (0.020–0.025).
-These preliminary results reveal no clear relationship between test MSE and the number of survey questions, suggesting that wave-specific factors, such as survey content, attribute coverage, or question formulation, may play a larger role in shaping performance.
-Figure 19: Test performance across ATP waves.
-Each point represents a wave-specific run, colored by wave identity.
-The horizontal dashed lines mark two threshold levels: 0.015 and 0.020, used to classify performance regimes.
-To illustrate the interaction between question-level entropy dynamics and predictive performance, we include three snapshots of the panel study in Figure 20, showcasing the differences in question entropy trajectories during active endowment generation and prediction performance.
-Figure 20: Representative snapshots for three ATP waves with different test-time MSE.
-Top: the worst-performing wave.
-Middle: a moderately performing wave.
-Bottom: the best-performing wave.
-Left: entropy trajectories classified as rising (green), falling (red), or stable (gray).
-Center/right: Lasso regression fits for aggregate response prediction.
-While rising entropy often coincides with improved fit, some waves achieve strong performance with limited entropy change, suggesting additional contributing factors beyond diversity expansion.
-Interestingly, the wave with the lowest test MSE (bottom row) exhibits relatively modest entropy change, while the poorest-performing wave (top row) shows mostly flat entropy evolutions.
-This suggests that while entropy dynamics can be indicative of diversity expansion, strong initial endowment coverage or other wave-specific factors (e.g., question content or noise) may also influence alignment outcomes.
-Entropy trends are a useful diagnostic but should be interpreted in the context of wave-specific factors.
-The patterns observed in the panel study warrant deeper analysis, which is likely to inform refinements that make P2P more stable and effective across settings.
-One especially promising direction is to extend the entropy-based active endowment generation strategy to support multi-objective criteria, allowing the system to balance diversity, coverage, and alignment more systematically.
-Another promising extension is to include a module that directly measures question semantics.
-Based on our preference reconstruction theory, survey questions define the response space $\mathcal{R}$.
-Question entropy serves as a key signal for underlying diversity when conditioning on a fixed question.
-However, an equally important source of variation lies in the diversity across questions themselves.
-This includes differences in topic coverage, difficulty, and framing—factors that shape the expressiveness and completeness of the response space.
-Ideally, the question set should span a semantically broad topic space so that variations in response entropy meaningfully reflect differences in latent preferences.
-A dedicated semantic module would enable content-aware data-splitting strategies and support more rigorous generalization analysis across question types.
-This is an important direction for future development of P2P.
-Due to budget constraints, our empirical evaluation relies on American Trends Panel (ATP) waves, which are topical in nature and tend to focus on a narrow set of issues per wave.
-While this allows for controlled testing, it limits the diversity of question types and themes.
-A more expansive and promising application domain for P2P is large-scale general social surveys, such as the General Social Survey (GSS) or World Values Survey, which span a wide range of topics and include high-quality, theoretically grounded questions relevant across disciplines.
-Testing P2P in such settings would provide a more comprehensive benchmark for its alignment and coverage capabilities, and yield better insights into its generalizability across domains of social inquiry.
-Moreover, the agent ensembles generated from general social surveys could, in turn, be used to provide feedback on smaller topical surveys to improve survey design and assess content quality.
-Appendix C Methodological and Experimental Details
-C.1 Prompts
-C.1.1 Attribute Learner (AttributeLearner)
-To generate attributes from a single question:
-You are an intelligent research assistant trained to analyze individual survey questions and infer which human attributes might influence how different people respond.
-You are given a single survey question.
-Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to shape responses to this question.
-Focus on underlying factors that would cause meaningful variation in answers across different types of people. Avoid generic or overly broad attributes.
-Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
-**Example output format:** ["religious affiliation", "political ideology", "trust in government"]
-To generate attributes from a set of questions:
-You are an intelligent research assistant trained to analyze survey questions and infer which human attributes might influence how individuals respond.
-You are given a set of training-only survey questions.
-Your task is to propose a list of relevant human attributes—such as demographics, beliefs, values, personality traits, or ideological leanings—that are likely to affect responses to these questions.
-Carefully analyze the content and framing of the questions.
-Identify underlying factors that might shape how different people respond.
-Focus on attributes that are salient, discriminative, and potentially variable across respondents.
-Respond **only** with a Python-style list of double-quoted strings. Do not include any explanation, headers, or prose before or after the list.
-**Example output format:** ["attribute1", "attribute2", "attribute3", "attribute4"]
-C.1.2 Endowment Model (EndowmentModel)
-To generate endowments from attributes:
-You are an expert assistant trained to generate realistic, diverse, and demographically plausible personas for social science surveys.
-Each persona should include:
-‘eid‘: a short, lowercase, variable-safe identifier that encodes key traits (e.g., urban_liberal_30s_female). No punctuation or spaces.
-‘endow_text‘: a natural language description of the persona (1-2 sentences), written as if describing a survey respondent.
-Instructions:
-Represent a wide range of age, gender, race, education, region, and political ideology.
-Avoid repetition of phrasing or demographic combinations across personas.
-Do not include explanations or formatting outside of the JSON array.
-C.1.3 Survey Conductor (SurveyConductor)
-To use an agent to answer a survey question:
-You are completing a survey.
-Your answer should reflect the person described in the profile above, using their preferneces, beliefs and experiences.
+Appendix D Empirical Experiments: Supplementary Results
+D.1 Main Results: Baseline Comparison on ATP Waves
+In our setup, MSE is computed on the option-level probabilities after binarizing each multiple-choice item, so MSE penalizes misallocation of probability mass across the full categorical distribution.
+To facilitate comparisons across works and corroborate the use of MSE as a valid distribution metric, we follow previous studies (Cao et al., 2025; Feng et al., 2024; Durmus et al., 2024) and additionally report Jensen-Shannon Divergence (1-JSD) and Earth Mover Distance (EMD) for the key results in this paper.
+Table 3 presents the full distribution metrics comparison of baselines and P2P across 14 ATP waves.
+P2P consistently outperforms Vanilla and PERSONA across all three metrics, registering lower MSE (0.014 vs.
+0.026–0.029), higher 1-JSD (0.828 vs.
+0.757–0.763), and lower EMD (0.067 vs.
+0.144–0.178).
+The 53% reduction in EMD is particularly notable, indicating that P2P predictions are structurally closer to ground-truth distributions.
+Table 3: Full distribution metrics comparison across 14 ATP waves.
+MSE measures mean squared error of predicted response probabilities.
+1-JSD (1 minus Jensen-Shannon Divergence) measures distributional similarity (higher is better).
+EMD (Earth Mover’s Distance) measures distributional distance (lower is better).
+P2P reports mean $\pm$ std over 3 runs.
+Vanilla PERSONA P2P
+Wave MSE 1-JSD EMD MSE 1-JSD EMD MSE 1-JSD EMD
+W26 .015 .817 .075 .021 .782 .078 $010_{\pm.000} .$868±.001 .056±.001
+W27 .030 .766 .419 .042 .730 .556 $022_{\pm.001} .$823±.006 .072±.003
+W29 .030 .757 .096 .030 .743 .107 $016_{\pm.005} .$815±.027 .083±.013
+W32 .020 .803 .310 .026 .768 .454 $016_{\pm.002} .$833±.007 .066±.003
+W34 .022 .763 .091 .020 .779 .086 $011_{\pm.001} .$822±.010 .070±.007
+W36 .033 .750 .087 .035 .761 .081 $019_{\pm.004} .$818±.013 .063±.006
+W41 .022 .759 .098 .023 .777 .092 $013_{\pm.001} .$828±.002 .069±.006
+W42 .034 .758 .086 .038 .750 .091 $011_{\pm.000} .$853±.003 .053±.001
+W45 .016 .804 .275 .028 .778 .482 $008_{\pm.002} .$848±.013 .057±.010
+W49 .045 .731 .095 .047 .731 .090 $021_{\pm.004} .$814±.017 .065±.005
+W50 .027 .708 .103 .024 .711 .111 $013_{\pm.004} .$805±.015 .076±.008
+W54 .022 .734 .092 .026 .740 .083 $014_{\pm.002} .$801±.019 .068±.007
+W82 .025 .742 .096 .019 .778 .079 $008_{\pm.002} .$856±.012 .052±.005
+W92 .025 .784 .093 .031 .775 .096 $017_{\pm.003} .$810±.008 .083±.004
+Avg .026 .763 .144 .029 .757 .178 $\mathbf{.014}$ .828 .067
+Table 4 shows stability analysis across four representative waves spanning weak, medium and strong alignment difficulty (characterized by test MSE).
+P2P ranks top on entropy, MSE, 1-JSD and EMD across all waves, with cost consistently between Vanilla and PERSONA.
+Baseline variance is negligible for MSE (max std = 0.005), entropy (max std = 0.007), 1-JSD (max std = 0.012), and cost (max std = 0.007).
+EMD shows higher baseline variance on W45 (std up to 0.227 for PERSONA, 0.125 for Vanilla), indicating occasional distributional instability that P2P avoids (max std = 0.010).
+The performance gap substantially exceeds error bands in most cases, confirming that improvements are statistically meaningful.
+We report 3 runs for P2P to verify robustness of the active endowment generation process, and single runs for baselines given their low variance.
+Table 4: Stability analysis on representative ATP waves.
+We select waves representing weak (W27), medium (W32), and strong (W42, W45) alignment performance.
+All values are mean $\pm$ std over 3 runs.
+Standard deviations are consistently small across all metrics, justifying single-run reporting for baselines in Table 1.
+Wave Method Entropy $\uparrow$ MSE $\downarrow$ 1-JSD $\uparrow$ EMD $\downarrow$ Cost (\$) $\downarrow$
+W27 (Weak) Vanilla .462 $\pm$ .007 .032 $\pm$ .005 .795 $\pm$ .008 .409 $\pm$ .019 0.63 $\pm$ .004
+PERSONA .402 $\pm$ .000 .041 $\pm$ .001 .728 $\pm$ .004 .557 $\pm$ .011 1.26 $\pm$ .000
+P2P .536 $\pm$ .010 .022 $\pm$ .001 .823 $\pm$ .006 .071 $\pm$ .003 0.71 $\pm$ .004
+W32 (Medium) Vanilla .551 $\pm$ .006 .019 $\pm$ .001 .801 $\pm$ .003 .350 $\pm$ .001 0.64 $\pm$ .004
+PERSONA .557 $\pm$ .000 .025 $\pm$ .000 .772 $\pm$ .004 .444 $\pm$ .008 1.28 $\pm$ .000
+P2P .618 $\pm$ .015 .016 $\pm$ .002 .833 $\pm$ .007 .066 $\pm$ .003 0.72 $\pm$ .001
+W42 (Strong) Vanilla .317 $\pm$ .007 .029 $\pm$ .005 .768 $\pm$ .010 .081 $\pm$ .004 0.86 $\pm$ .007
+PERSONA .287 $\pm$ .000 .036 $\pm$ .001 .753 $\pm$ .003 .090 $\pm$ .001 1.71 $\pm$ .000
+P2P .445 $\pm$ .026 .010 $\pm$ .002 .863 $\pm$ .019 .048 $\pm$ .007 0.97 $\pm$ .004
+W45 (Strong) Vanilla .389 $\pm$ .002 .015 $\pm$ .003 .805 $\pm$ .012 .146 $\pm$ .125 0.62 $\pm$ .004
+PERSONA .365 $\pm$ .001 .029 $\pm$ .001 .773 $\pm$ .006 .219 $\pm$ .227 1.24 $\pm$ .000
+P2P .445 $\pm$ .006 .008 $\pm$ .002 .848 $\pm$ .013 .057 $\pm$ .010 0.69 $\pm$ .002
+Table 5 compares lasso and elastic net as Stage 2 regression choices for P2P.
+Performance is nearly identical across all waves (average $\Delta$ MSE $\approx$ 0.0001, with 10 of 14 waves showing no difference).
+As lasso typically selects a more compact ensemble than elastic net, we prioritize its use as the default regression method.
+Table 5: Comparison of lasso and elastic net regression for P2P across 14 ATP waves.
+Values are mean $\pm$ std over 3 runs.
+Both methods yield nearly identical performance, demonstrating robustness to regularization choice.
+Wave Lasso MSE ENet MSE $\Delta$ Winner
+W26 .0098 $\pm$ .0002 .0098 $\pm$ .0002 .0000 Tie
+W27 .0218 $\pm$ .0010 .0238 $\pm$ .0027 .0020 Lasso
+W29 .0156 $\pm$ .0049 .0156 $\pm$ .0049 .0000 Tie
+W32 .0160 $\pm$ .0019 .0160 $\pm$ .0019 .0000 Tie
+W34 .0111 $\pm$ .0011 .0111 $\pm$ .0011 .0000 Tie
+W36 .0191 $\pm$ .0042 .0187 $\pm$ .0037 .0004 ENet
+W41 .0126 $\pm$ .0009 .0126 $\pm$ .0009 .0000 Tie
+W42 .0110 $\pm$ .0003 .0101 $\pm$ .0006 .0009 ENet
+W45 .0079 $\pm$ .0023 .0079 $\pm$ .0023 .0000 Tie
+W49 .0211 $\pm$ .0037 .0211 $\pm$ .0037 .0000 Tie
+W50 .0133 $\pm$ .0039 .0133 $\pm$ .0039 .0000 Tie
+W54 .0144 $\pm$ .0018 .0146 $\pm$ .0018 .0002 Lasso
+W82 .0083 $\pm$ .0016 .0083 $\pm$ .0016 .0000 Tie
+W92 .0168 $\pm$ .0029 .0173 $\pm$ .0023 .0005 Lasso
+Avg .0142 .0143 .0001
+D.1.1 Clustering
+Based on Vanilla and PERSONA, we design two variants using K-means clustering as a pre-processing step.
+Specifically, we first expand the initial pool of personas for Vanilla and PERSONA to 1000, computing their embeddings (via OpenAI’s text-embedding-3-large) and cluster them (k=300).
+We pick the personas (medoids) closest to the centroids to construct the candidate pool and then run our regression stage on the resulting pool.
+Table 6 shows results on ATP W42.
+Table 6: Comparison of Vanilla, PERSONA and their clustering-based variants where 1000 agents are reduced to 300 via $k$-means clustering.
+Metric Vanilla Vanilla PERSONA PERSONA
+[…]
+Avg. Entropy $\uparrow$ 0.3172 $\pm$ 0.0069 0.3162 0.2871 $\pm$ 0.00026 0.2844
+Test MSE (Lasso) $\downarrow$ 0.0285 $\pm$ 0.00490 0.0266 0.0362 $\pm$ 0.00116 0.0395
+Test MSE (Enet) $\downarrow$ 0.0286 $\pm$ 0.00476 0.0266 0.0339 $\pm$ 0.00124 0.0387
+Cost (USD) $\downarrow$ 0.8649 $\pm$ 0.0070 0.8834 1.7129 $\pm$ 0.000016* 1.7203*†
+* Only considering survey elicitation cost on 300 personas—actual cost is higher.
+Excluding embedding cost.
+Clustering yields a test MSE of 0.0266 for Vanilla that lies well within the variability band of the original Vanilla baseline (0.0285 $\pm$ 0.0049) and slightly worse MSE for PERSONA (0.0395 vs.
+0.0362 $\pm$ 0.0011), with no conspicuous change on entropy and a slight increase in cost (for Vanilla due to extra generation).
+Hence, clustering does not yield a stronger baseline.
+This matches our intuition for two reasons.
+First, when we apply clustering on top of the existing baselines (starting from the same 300 endowments/personas), it can only discard information that the regression step could have used.
+In other words, it weakens the candidate pool without giving L1 regression any new capability.
+Second, even in the more lenient setting where we give clustering a head start by generating 1000 candidates and then compressing them to 300, the mechanism still does not produce a stronger baseline.
+As clustering operates in the embedding space and tends to group personas by textual or stylistic similarity, it does not necessarily track preference-relevant variation in survey responses.
+As a result, minority or low-density preference types are prone to being averaged away in embedding clustering, whereas our regression layer selects proxies directly based on their contribution to matching the observed response distributions.
+This corroborates the design of our main baselines.
+D.1.2 Raw Model Performance
+To better contextualize the gain of the ensemble methods (Vanilla, PERSONA, and P2P), we provide the raw model backend performance as a reference baseline.
+To construct the baseline, we directly query the model backend (gemini-2.0-flash) for answers to the survey questions, using the following instruction:
+You are completing a public opinion survey.
+Answer each question as an average adult respondent in the US population, based only on the question text and options.
 Respond with only the final answer string, not the code or label in brackets.
@@
 Just return the answer text exactly as it appears in the options.
-C.2 Active Endowment Generation
-Algorithm 1 Active Endowment Generation
-1: Initialize attribute bank $\mathcal{B}$ with core, thematic, and theoretical templates
-2: Derive survey-specific attributes from $\mathcal{Q}_{\mathrm{train}}$ via AttributeLearner; append to $\mathcal{B}$ Initial Sampling Stage
-3: Sample initial endowments $\mathcal{E}_{0}$ via equal-mode sampling from $\mathcal{B}$
-4: Instantiate initial agents $\mathcal{A}_{0}$ from $\mathcal{E}_{0}$ and elicit responses $\mathcal{R}_{0}$ on $\mathcal{Q}_{\mathrm{train}}$
-5: Initialize ThemeVariabilityTracker (henceforce, Tracker) with $\mathcal{A}_{0}$, $\mathcal{R}_{0}$, and $\mathcal{Q}_{\mathrm{train}}$. Expansion Stage
-6:while agent budget $N_{A}$ not yet reached do
-7:  Compute variability scores and mode sampling probabilities via Tracker
-8:  Allocate endowment budget across modes according to probabilities
-9:  if question patching is enabled then
-10:   Identify lowest-entropy questions $\mathcal{Q}_{\mathrm{low}}$ via Tracker
-11:   for each $q\in\mathcal{Q}_{\mathrm{low}}$ do
-12:    if $q$ is queried for the first time then
-13:     Extract attributes for $q$ via AttributeLearner and append to $\mathcal{B}$
-14:    else if $q$ has appeared more than $n$ times then
-15:     Enable mixed-mode strategy using $q$ and top-performing mode
-16:    end if
-17:   end for
-18:  end if
-19:  Generate new endowments $\mathcal{E}_{\mathrm{new}}$ using allocated budget over sampled modes, patched question modes (if any) and mixed modes (if any).
-20:  Instantiate agents $\mathcal{A}_{\mathrm{new}}$ and elicit responses $\mathcal{R}_{\mathrm{new}}$ on $\mathcal{Q}_{\mathrm{train}}$
-21:  Update: $\mathcal{A}\leftarrow\mathcal{A}\cup\mathcal{A}_{\mathrm{new}}$, $\mathcal{R}_{\mathcal{A}}\leftarrow\mathcal{R}_{\mathcal{A}}\cup\mathcal{R}_{\mathrm{new}}$
-22:  Update Tracker with $\mathcal{A}$, $\mathcal{R}_{\mathcal{A}}$, and $\mathcal{Q}_{\mathrm{train}}$
-23:end while
-24:Return: Final agent pool $\mathcal{A}$ and full response set $\mathcal{R}_{\mathcal{A}}$
-Algorithm 2 Tracker Update Procedure
-1: Input: Current agent pool $\mathcal{A}$, current agent responses $\mathcal{R}_{\mathcal{A}}$, and training questions $\mathcal{Q}_{\mathrm{train}}$
-2:for each question $i\in\mathcal{Q}_{\mathrm{train}}$ do
-3:  Compute question entropy $H_{i}(\mathcal{A})$
-4:end for
-5:for each mode do
-6:  Compute variability score $V(\mathcal{A}_{\mathrm{mode}})$
-7:end for
-8: Compute softmax probabilities from variability scores
-9: Output: Variability scores, question entropies, and sampling probabilities.
-C.3 Example Endowments
-An example of an endowment for a progressive female in her 20s living in an urban area
+We draw 300 Monte-Carlo samples per question with temperature set to 0.7.
+This results in an empirical answer distribution per question.
+On ATP W42, this baseline attains test MSE 0.0918 and average entropy 0.0161, conspicuously worse than the results for the ensemble baselines, confirming the gains via ensemble methods (Vanilla, PERSONA, and P2P).
+D.2 Locale Generalization: World Values Survey
+The main paper results focus on the American Trends Panel (ATP), which represents a single survey program targeting the US population.
+While our 14-wave study demonstrates that P2P generalizes across multiple topics and different cross sections, it remains intriguing how it performs at other locales and on different survey programs.
+To address this question, we evaluate P2P on the World Values Survey (WVS) Wave 7 (Haerpfer et al., 2022), using cleaned data from Cao et al.
+(2025).
+We select three locales representing different cultural and linguistic contexts: United States, as it provides a bridge to our ATP findings, enabling direct comparison within the same population under a different survey program; Great Britain, which represents Western Europe, sharing linguistic and value overlap with the US but differing in political and cultural contexts; and Hong Kong SAR, a multicultural hub in East Asia, allowing us to test on populations less represented in typical LLM training data.
+D.2.1 Experiment Setup
+For the WVS experiment, we use a similar setup to the ATP study: survey questions are split randomly according to a $7:1.5:1.5$ ratio into the training, validation, and test sets; endowment budget set is set at 300; initial sampling generates $10$ endowments for each mode; active endowment generation is run in 5 update steps with the preset attribute bank, question patching (lowest 3; 75% endowment budget); 10 attributes drawn for each endowment generation; Gemini-2.0-flash is used as the backend.
+Different from the ATP study, we additionally pass the locale information (US, GB, HK) to the endowment model to ensure regional relevance of generated personas.
 […]
-A 24-year-old urban female who identifies as bisexual and holds a master’s degree in social work.
-She exhibits a high fairness bias and is often concerned about equality.
-Despite her awareness, she sometimes struggles with extreme response bias, influenced by her strong convictions.
-An example of an endowment for a nonbinary in their 30s valuing liberty living in a suburban area
-[…]
-They are a 32-year-old nonbinary software developer living in the suburbs.
-They highly value personal liberty and have a strong work ethic, often putting in extra hours for their projects.
-With a master’s degree in computer science, they frequently engage with scientific research, feeling that understanding complex topics fosters innovation.
-Their income is above average, and they embrace intelligence as a vital asset for personal growth.
-An example of an endowment for a white male progressive in his 60s living in a small town
-[…]
-Pat is a 67-year-old White male from a small town.
-He identifies as a progressive and believes strongly in the importance of intelligence in community discussions and personal growth.
-Pat, who identifies as heterosexual, is skeptical of the increasing influence of big corporations on local politics, and he often seeks media that highlights grassroots movements and community activism.
-He balances his media consumption between traditional outlets and local storytelling platforms, knowing the value of good journalism.
-Appendix D Extended Literature Review
-D.1 Large Language Models to Emulate Human Preferences
-Leveraging the ability of LLMs to emulate human or human-group responses with considerable fidelity, LLM alignment or personalization Tseng et al.
-[2024] has been largely focused on emulating human responses using various techniques.
-These include learning from feedback or reward models Bai et al.
-[2022], Christiano et al.
-[2017], under distribution shifts Leike et al.
-[2018], Krueger et al.
-[2021], finetuning Tan et al.
-[2024], retrieval augmented generation Prahlad et al.
-[2025], representation learning Ren et al.
-[2024], and using human-generated Horton [2023], Castricato et al.
-[2025], Argyle et al.
-[2023] prompts.
+Table 7 reports results across the three locales. Each experiment is repeated 3 times.
+D.2.2 Results and Discussion
+Table 7: P2P performance on WVS Wave 7 (US, Great Britain, Hong Kong SAR) under a 7:1.5:1.5 question split, using gemini-2.0-flash as the backend. Reported are mean $\pm$ std over 3 repeated runs.
+Metric WVS W7 US WVS W7 GB WVS W7 HK
+Avg. Entropy $\uparrow$ 0.5720 $\pm$ 0.0149 0.5710 $\pm$ 0.0044 0.5478 $\pm$ 0.0211
+Cost (USD) $\downarrow$ 1.7551 $\pm$ 0.0097 1.7095 $\pm$ 0.0031 1.7365 $\pm$ 0.0080
+Test MSE (Lasso) $\downarrow$ 0.0149 $\pm$ 0.0005 0.0141 $\pm$ 0.0014 0.0242 $\pm$ 0.0013
+Test 1-JSD (Lasso) $\uparrow$ 0.8030 $\pm$ 0.0055 0.7895 $\pm$ 0.0049 0.7590 $\pm$ 0.0046
+Test MCPA (Lasso) $\uparrow$ 0.6667 $\pm$ 0.0152 0.6481 $\pm$ 0.0321 0.5789 $\pm$ 0.0263
+Test EMD (Lasso) $\downarrow$ 0.1075 $\pm$ 0.0035 0.1038 $\pm$ 0.0081 0.1424 $\pm$ 0.0034
+Test MSE (Elastic Net) $\downarrow$ 0.0147 $\pm$ 0.0005 0.0137 $\pm$ 0.0010 0.0244 $\pm$ 0.0005
+Test 1-JSD (Elastic Net) $\uparrow$ 0.8066 $\pm$ 0.0067 0.7950 $\pm$ 0.0073 0.7602 $\pm$ 0.0032
+Test MCPA (Elastic Net) $\uparrow$ 0.6930 $\pm$ 0.0152 0.6296 $\pm$ 0.0321 0.5877 $\pm$ 0.0304
+Test EMD (Elastic Net) $\downarrow$ 0.1066 $\pm$ 0.0025 0.1004 $\pm$ 0.0054 0.1430 $\pm$ 0.0046
+Table 8: Comparison of P2P performance on ATP (US) and WVS (US). ATP values are mean $\pm$ std over 14 waves; WVS values are mean $\pm$ std over 3 runs.
+Metric ATP (14 waves) WVS US
+Test MSE (Lasso) $\downarrow$ 0.014 $\pm$ 0.004 0.015 $\pm$ 0.001
+Avg. Entropy $\uparrow$ 0.53 $\pm$ 0.06 0.57 $\pm$ 0.01
+Cost (USD) $\downarrow$ 0.8 $\pm$ 0.2 1.76 $\pm$ 0.01
+As shown in Table 7, performance is higher for US (MSE $\approx$ 0.015) and GB (MSE $\approx$ 0.014) and somewhat lower for HK (MSE $\approx$ 0.024), aligning with broader findings that large language models trained primarily on English and Western-centric data tend to better match Western public opinion than non-Western populations (Durmus et al., 2024).
+The HK performance gap motivates the error analysis presented in Section 4.4 of the main paper.
+Despite the gap, the HK results remain within the performance range observed across our 14-wave ATP study presented in the main paper (cf.
+Table 1), suggesting that while locale affects difficulty, P2P remains broadly applicable.
+Compared with ATP, WVS results register a higher per-survey cost at approximately \$1.73, due to the higher number of survey questions (185 for US and GB, 172 for HK).
+Notably, the US results are nearly identical across survey programs (ATP: 0.014; WVS: 0.015), providing evidence against program-specific overfitting (Table 8).
+D.3 Ablation Studies
+D.3.1 Attributes
+As discussed in Section 3.2, attributes play the role of control handles in P2P’s generation logic, and freeform templates allow P2P to derive attributes from a specific survey or question, complementing the preset attribute bank with data-driven insights on endowment generation.
+To evaluation the potency of the freeform templates, we conduct an ablation in study in which we sequentially turn off question patching, survey patching and both.
+We additionally test the impact of per-endowment attribute cap through a setting where the default max attribute=10 is changed to 20.
+Table 9: Attribute Ablation. Reported are mean $\pm$ std over 3 repeated runs.
+Metric Avg. Entropy $\uparrow$ Test MSE $\downarrow$ Test MSE $\downarrow$ Cost (USD) $\downarrow$
+(Lasso) (Enet)
+Original 0.4451 $\pm$ .0256 0.0095 $\pm$ .0024 0.0079 $\pm$ .0026 0.9708 $\pm$ .0036
+Survey patch off 0.4378 $\pm$ .0134 0.0103 $\pm$ .0011 0.0102 $\pm$ .0002 0.9785 $\pm$ .0122
+Question patch off 0.3989 $\pm$ .0203 0.0114 $\pm$ .0016 0.0109 $\pm$ .0015 0.9698 $\pm$ .0011
+Freeform off 0.3953 $\pm$ .0097 0.0128 $\pm$ .0008 0.0123 $\pm$ .0012 0.9688 $\pm$ .0015
+max attributes = 20 0.4363 $\pm$ .0061 0.0124 $\pm$ .0046 0.0115 $\pm$ .0041 0.9880 $\pm$ .0167
+Results in Table 9 suggest that performance of P2P is robust to attribute cap, whereas turning off freeform templates (especially question patching) degrades both average entropy and test MSE without significant cost gain.
+Therefore, from a performance vantage, it is recommended to keep the freeform templates for the extra benefits of data-driven insights.
+D.3.2 Endowment Budget
+We run P2P for 9 different endowment budgets, scaled linearly from 130 to 450.
+We increase the update steps of active endowment generation accordingly to ensure that at each step the number of new endowments generated is fixed at 20.
+We also include a budget of 110 with no updates as an ablation of adaptive sampling.
+Table 10 shows that on W42, a higher budget generally improve test performance but gains plateau after about three updates, indicating a moderate budget suffices.
+Table 10: Performance of P2P with varying endowment budget.
+The average question entropy, test MSE (Lasso and ElasticNet), and generation cost are repeated.
+All values are mean $\pm$ std over 3 repeated runs.
+(a) Budgets 110–210
+Endowment Budget 110 130 170 210
+Avg. Entropy $\uparrow$ 0.3777 $\pm$ 0.0191 0.3697 $\pm$ 0.0057 0.4180 $\pm$ 0.0080 0.4341 $\pm$ 0.0043
+Test MSE (Lasso) $\downarrow$ 0.0144 $\pm$ 0.0018 0.0124 $\pm$ 0.0043 0.0167 $\pm$ 0.0041 0.0100 $\pm$ 0.0025
+Test MSE (Enet) $\downarrow$ 0.0149 $\pm$ 0.0014 0.0123 $\pm$ 0.0045 0.0156 $\pm$ 0.0028 0.0103 $\pm$ 0.0026
+Cost (USD) $\downarrow$ 0.3487 $\pm$ 0.0035 0.8043 $\pm$ 0.0032 0.4108 $\pm$ 0.0089 0.5453 $\pm$ 0.0063
+(b) Budgets 250–330
+Endowment Budget 250 290 330
+Avg. Entropy $\uparrow$ 0.4179 $\pm$ 0.0165 0.4369 $\pm$ 0.0168 0.4308 $\pm$ 0.0208
+Test MSE (Lasso) $\downarrow$ 0.0111 $\pm$ 0.0006 0.0090 $\pm$ 0.0013 0.0133 $\pm$ 0.0030
+Test MSE (Enet) $\downarrow$ 0.0104 $\pm$ 0.0006 0.0097 $\pm$ 0.0017 0.0121 $\pm$ 0.0031
+Cost (USD) $\downarrow$ 0.6715 $\pm$ 0.0129 0.9365 $\pm$ 0.0027 1.0688 $\pm$ 0.0119
+(c) Budgets 370–450
+Endowment Budget 370 410 450
+Avg. Entropy $\uparrow$ 0.4216 $\pm$ 0.0179 0.4540 $\pm$ 0.0142 0.4433 $\pm$ 0.0092
+Test MSE (Lasso) $\downarrow$ 0.0140 $\pm$ 0.0037 0.0080 $\pm$ 0.0006 0.0099 $\pm$ 0.0020
+Test MSE (Enet) $\downarrow$ 0.0136 $\pm$ 0.0055 0.0084 $\pm$ 0.0004 0.0097 $\pm$ 0.0020
+Cost (USD) $\downarrow$ 1.2050 $\pm$ 0.0179 1.3295 $\pm$ 0.0142 1.4619 $\pm$ 0.0092
+D.3.3 Model Backend
+We test P2P’s performance on five model backends: GPT-4.1-mini, GPT-4.1-nano, Gemini-2.0-flash, Gemini-2.5-flash and a locally hosted Qwen.
+Performance varies substantially, with the best average question entropy (0.471) and test MSE (0.009) achieved by GPT-4.1-mini (Table 11).
+Gemini-2.0-flash ranks second on these metrics but is markedly more cost-efficient, requiring only one quarter of the cost of GPT-4.1-mini.
+Smaller models, like GPT-4.1-nano and Gemini-2.5-flash-lite, perform worse, consistent with our theoretical intuition that limited model capacity constrains the learnability condition (Definition 2.2).
+Table 11: Comparison of P2P performance on W42 using different model backends.
+Reported are average question entropy, test MSE (Lasso and ElasticNet), and generation cost.
+For each model, we fix the endowment budget at 300.
+All values are mean $\pm$ std over 3 repeated runs.
+Model Avg. Entropy Test MSE (Lasso) Test MSE (Enet) Cost (USD)
+GPT-4.1-mini 0.471 0.0090 0.0088 3.83
+(0.017) (0.0014) (0.0014) (0.045)
+GPT-4.1-nano 0.3996 0.0226 0.0213 0.9265
+(0.0154) (0.0042) (0.0048) (0.0043)
+Qwen 0.193 0.0204 0.0196
+(0.040) (0.0014) (0.0027)
+Gemini-2.0-flash 0.445 0.0104 0.0091 0.971
+(0.026) (0.0038) (0.0046) (0.004)
+Gemini-2.5-flash-lite 0.360 0.0149 0.0151 1.023
+(0.0008) (0.0009) (0.0013) (0.0010)
+Note. Qwen was run on a local server, thereby incurring no API cost.
+D.3.4 Regression-based aggregation
+To gauge the contribution of the regression-based aggregation module, we ablate it by replacing regression with simple averaging of agent responses, which increases the test MSE from 0.0104 to 0.0254 (Table 12).
+In essence, regression not only distills the agent collection into a more compact ensemble, but it also plays a crucial role in ensuring the population preference reflected in the survey data are actually learned by the ensemble.
+Table 12: Comparison of P2P performance with and without the regression-based aggregation stage on ATP W42.
+In the ablation, agent responses from active endowment generation are combined by simple averaging.
+For the regression setting, we use test MSE for Lasso.
+Metric Without regression With regression
+Test MSE (Lasso) $\downarrow$ 0.0254 $\pm$ 0.00041 0.0104 $\pm$ 0.00383
+We also compare the default option-level weighting against block weighting, where each question contributes equally to the loss. Performance remains comparable under both schemes for W42 (Table 13).
+Table 13: Sensitivity check of weighting schemes. Reported are mean $\pm$ std over 3 repeated runs.
+Weighting Train MSE $\downarrow$ Test MSE $\downarrow$ Train MSE $\downarrow$ Test MSE $\downarrow$
+Scheme (Lasso) (Lasso) (Enet) (Enet)
+Default 0.00526 $\pm$ .00141 0.00946 $\pm$ .00241 0.00353 $\pm$ .00161 0.00792 $\pm$ .00261
+Block 0.00472 $\pm$ .00228 0.00884 $\pm$ .00281 0.00389 $\pm$ .00152 0.00797 $\pm$ .00263
+D.4 Stress Test: Comparison with an SFT-Aligned Model under Topic Shift
+P2P is designed as a plug-and-play preference alignment method that can be deployed on new populations with limited-data and limited-compute regimes.
+In the main paper, the baseline comparisons thus focus on prompting-based ensemble methods under the same constraints.
+As discussed in Appendix E, supervised finetuning (SFT) for a distribution-calibrated model offers an alternative approach under more lenient data and compute environments.
+An exemplar is Cao et al.
+(2025), where SFT is used to train a predictive model based on pooled World Values Survey (WVS) data from 46 countries, and conditioned on locale to directly predict answer distribution for a given survey question.
+They further test the SFT-aligned model on a distinct unseen set of WVS questions along with other cross-country and cross-dataset validation tasks.
+In this regard, our work differs from Cao et al.
+(2025) not just in terms of training regimes, but also in key methodological premises.
+Specifically, we regard each survey panel as a benchmark for a specific population preference on a given topic at a given moment in time, where such preference is gaugeable through survey responses (cf.
+Figure 1, right panel).
+We do not assume that preferences on a topic domain can be reliably inferred if the training questions never probe that domain.
+For example, inferring political attitudes solely from non-political or religious items is a strong assumption for both humans and synthetic agents.
+Likewise, we do not assume preferences can carry over from one survey program to another, particularly when the surveys are conducted at different points in time and on different subpopulations of the same locale.
+Studying how aligned preferences generalize across topics, survey programs and over time is itself a key direction for future work, and will require both theoretical and empirical advances.
+Nevertheless, Cao et al.
+(2025)’s setup constitutes an interesting stress test for P2P outside our studied scenarios.
+We present in this section a head-to-head comparison with their method on their WVS arena.
+We follow their protocol: Q1 and Q2 for train/validation and Q3 as test, which induces a topic shift from general attitudinal & religious/ethical items (Q1+Q2) to political interest and culture items (Q3).
+In our language, this asks: if we align to a population on broad attitudinal dimensions (Q1+Q2), how well does the aligned ensemble generalize to a new unseen dimension (Q3: political-cultural questions)?
+As in Section D.2, we focus on the three representative locales: US (overlaps ATP), Great Britain (Western Europe) and Hong Kong SAR (East Asia).
+For each, we filter Cao et al.
+(2025)’s data to that locale, train P2P (backend: gemini-2.0-flash; 300 endowments) on local Q1+Q2 (train+val), and evaluate on Q3.
+Table 14 shows the training schemes for P2P and Cao et al.
+(2025).
+Table 14: Different training regimes for P2P and (Cao et al., 2025).
+Setup Train+Val Entries Test Entries Model backend
+(Q1 + Q2) (Q3)
+P2P (US) 185 (150 + 35) 59 gemini-2.0-flash
+P2P (GB) 172 (140 + 32) 60 gemini-2.0-flash
+P2P (HK) 185 (150 + 35) 59 gemini-2.0-flash
+Cao et al. (C1: 46 countries) 8427 (6841 + 1586) 2719 Llama3-8B-Instruct
+Table 15: P2P vs. Cao et al. (2025) on WVS (Q3). Reported are mean $\pm$ std over 3 repeated runs.
+Setup Method Test MSE $\downarrow$ Test 1-JSD $\uparrow$ Test MCPA $\uparrow$ Test EMD $\downarrow$
+P2P (US) Lasso 0.0217 $\pm$ .0014 0.7492 $\pm$ .0036 0.6271 $\pm$ .0143 0.1188 $\pm$ .0048
+P2P (US) Enet 0.0218 $\pm$ .0019 0.7497 $\pm$ .0051 0.6384 $\pm$ .0317 0.1184 $\pm$ .0066
+P2P (GB) Lasso 0.0249 $\pm$ .0009 0.7416 $\pm$ .0055 0.4889 $\pm$ .0347 0.1296 $\pm$ .0045
+P2P (GB) Enet 0.0249 $\pm$ .0009 0.7463 $\pm$ .0067 0.5111 $\pm$ .0096 0.1294 $\pm$ .0048
+P2P (HK) Lasso 0.0284 $\pm$ .0023 0.7115 $\pm$ .0125 0.4802 $\pm$ .0353 0.1507 $\pm$ .0122
+P2P (HK) Enet 0.0290 $\pm$ .0023 0.7151 $\pm$ .0133 0.4689 $\pm$ .0298 0.1532 $\pm$ .0108
+Cao et al. SFT 0.777 0.43 *
+* EMD not reported due to incompatible definition in Cao et al. (2025).
+Table 15 presents the results on the performance comparison.
+In summary, P2P achieves 1-JSD on Q3 comparable to Cao et al.
+(2025)’s SFT model and matches or exceeds their majority-class prediction accuracy (MCPA), despite using far fewer training entries per locale and no fine-tuning (Table 14).
+Following our earlier discussion, we interpret performance on Q3 as evidence that some preference structure transfers across topics, rather than as a requirement that a preference-alignment method must perfectly extrapolate to arbitrary unseen domains.
+Appendix E Extended Literature Review
+E.1 Pluralistic Alignment
+AI alignment refers to the process of ensuring that an AI system operates in accordance with human intentions and values, whether at an individual or aggregate level (Ji et al., 2023; Leike et al., 2018).
+As AI systems are adopted by increasingly diverse users, they must be designed to recognize and address a wider range of needs.
+This necessitates for pluralistic systems capable of capturing and representing the diversity of human values and perspectives (Sorensen et al., 2024b).
+Pluralistic alignment therefore seeks to align models not with a single gold standard, but with a diverse range of user preferences across various attributes.
+Most works on pluralistic alignment touch on three complementary ways, namely overton, steerable, and distributional pluralism, in which a single AI model or system can support diversity of views.
+In overton pluralism, the model aims to output a whole spectrum of reasonable responses (Sorensen et al., 2024a).
+Methods to achieve this include looking at alignment shifts to transform distributions toward overton coverage (Lake et al., 2024) or few shot prompting from community-specialized LLMs (Feng et al., 2024).
+In steerable pluralism, models are typically steered to adopt or favor particular perspectives or value attributes, often through few-shot examples (Feng et al., 2024; Adams et al., 2025).
+By contrast, our approach steers purely through attributes, with the goal of constructing diverse agents that embody different perspectives.
+In distributional pluralism, the distribution of the model over possible answers is intended to match that of some target population (Sorensen et al., 2024b), often achieved through model fine-tuning on human responses (Cao et al., 2025).
+Another way to categorize pluralistic alignment work is by the amount of data they rely on.
+While most approached in pluralistic alignment involve post-training, they range from resource-intensive approaches that fine-tune on large annotated corpora to lightweight methods that operate with minimal or no additional data.
+With sufficient data, specialized community models can be fine-tuned and subsequently combined using model merging techniques (Yuan et al., 2024), few-shot prompting from community model outputs (Feng et al., 2024), or federated averaging (Srewa et al., 2025) to scale with user diversity.
+If limited data are available, an alternative line of post-training approaches leverage in-context learning, where similarity-based retrieval (Adams et al., 2025) or group-informed retrieval (Chen et al., 2024) is performed to select few-shot examples that guide models toward pluralistic alignment.
+In the absence of detailed data from individual pluralistic users, a third post-training approach is inference-time conditioning.
+Inference-time conditioning uses structured prompts that encode sociodemographic traits or behavioral dispositions to condition an LLM to emulate the responses of a specific agent.
+These agents are also known as endowments, personas (Castricato et al., 2025) or Silicon Samples (Horton, 2023; Argyle et al., 2023) and are lightweight to deploy and data-efficient.
+As inference-time conditioning is cheap and fast, they enable mixture or vector-based formulations of pluralistic alignment, which combine agents via multi-objective rewards to avoid averaging out minority perspectives (Chen et al., 2025; Feng et al., 2025).
+Therefore, P2P follows this line of work for pluralistic alignment.
+In the first step, P2P performs inference-time conditioning to generate a sufficient set of agents that adequately span the preference space.
+Thereafter, in the second step, it combines agents to achieve a compact yet pluralistic set of preferences that represent the target survey population.
+E.1.1 Attributes
+Pluralistic alignment works that use inference-time conditioning often steer agents toward specific profiles using attributes that include moral and value dimensions (Adams et al., 2025), or broad morals, values, characteristics, and perspectives (Sorensen et al., 2024a).
+Attributes may also refer to demographic features, that are more tangible.
+Examples of these attributes include age, sex, education, income, and religion (Castricato et al., 2025).
+P2P uses the term attributes more broadly to encompass both moral and value dimensions and demographic features.
+Bundling these together allows P2P to define and condition LLM agents in a way that reflects both who they are and how they evaluate options.
+Other related works on pluralistic alignment relate to the curation of datasets, which support research on pluralistic alignment (Sorensen et al., 2024a; Zhou et al., 2025).
+E.2 Large Language Models to Emulate Human Preferences
+Leveraging the ability of LLMs to emulate human or human-group responses with considerable fidelity, LLM alignment or personalization (Tseng et al., 2024) has been largely focused on emulating human responses using various techniques.
+These include learning from feedback or reward models (Bai et al., 2022; Christiano et al., 2017), under distribution shifts (Leike et al., 2018; Krueger et al., 2021), finetuning (Tan et al., 2024), retrieval augmented generation (Prahlad et al., 2025), representation learning (Ren et al., 2024), and using human-generated (Horton, 2023; Castricato et al., 2025; Argyle et al., 2023) prompts.
 However, when diverse responses need to be generated from many users or user groups instead of individuals, the data and time resources required for alignment grow linearly with the number of LLMs that need to be aligned.
-To reduce alignment costs, some works instead rely on AI-generated Schuller et al.
-[2024], Salminen et al.
-[2024], Ge et al.
-[2025], Simmons [2023] prompts, or alignment to multiple user groups Zhao et al.
-[2024], Mondal et al.
-[2025], even entire populations Cao et al.
-[2025].
-However, even the best models struggle with this task Cao et al.
-[2025].
-Furthermore, responses from LLMs have been found to have less variation than responses from real surveys Bisbee et al.
-[2024], and LLMs have a tendency to respond in the middle category Wang et al.
-[2024].
+To reduce alignment costs, some works instead rely on AI-generated (Schuller et al., 2024; Salminen et al., 2024; Ge et al., 2025; Simmons, 2023) prompts, or alignment to multiple user groups (Mondal et al., 2025; Zhao et al., 2024b), even entire populations (Cao et al., 2025).
+However, even the best models struggle with this task (Cao et al., 2025).
+Furthermore, responses from LLMs have been found to have less variation than responses from real surveys (Bisbee et al., 2024), and LLMs have a tendency to respond in the middle category (Wang et al., 2024).
 Consequently, it is imperative to address these limitations and devise data- and time-efficient approaches that enable LLMs to more accurately and reliably serve as proxies for survey populations.
 Our work contributes to this goal by introducing an active learning–inspired alignment framework that dynamically constructs a diverse agent basis under resource constraints and reconstructs population-level preferences via regression-based aggregation, without relying on ground-truth demographic profiles.
-D.2 Data Selection and Active Learning
+E.3 Data Selection and Active Learning
 To increase preference coverage while operating under budget constraints, our active endowment generation pipeline draws inspiration from machine learning paradigms, particularly active learning.
-Traditional data selection methods aim to improve the data efficiency of supervised deep learning models by identifying the most informative or important training examples for generalization Coleman et al.
-[2020], Paul et al.
-[2021].
-Selection via Proxy (SVP) Coleman et al.
-[2020] uses a lightweight proxy model to estimate sample utility at a fraction of the cost, while GraNd and EL2N scores Paul et al.
-[2021] identify important training examples early in training.
-Similarly, deep active learning applies query strategies to iteratively select the most informative samples from a large pool of data to be added to the training dataset for retraining or fine-tuning Astorga et al.
-[2024], Hübotter et al.
-[2025], Li et al.
-[2025].
+Traditional data selection methods aim to improve the data efficiency of supervised deep learning models by identifying the most informative or important training examples for generalization (Coleman et al., 2020; Paul et al., 2021).
+Selection via Proxy (SVP) (Coleman et al., 2020) uses a lightweight proxy model to estimate sample utility at a fraction of the cost, while GraNd and EL2N scores (Paul et al., 2021) identify important training examples early in training.
+Similarly, deep active learning applies query strategies to iteratively select the most informative samples from a large pool of data to be added to the training dataset for retraining or fine-tuning (Astorga et al., 2024; Hübotter et al., 2025; Li et al., 2025).
 Active learning strategies balance exploitation, focusing on uncertain or high-impact samples, with exploration, seeking diverse and novel samples to broaden coverage.
@@
 Together, these components form an active learning–inspired loop that incrementally expands coverage of the preference space while minimizing redundant agent generation.
-Appendix E Mathematical Formalization of the Preference Reconstruction Theory
-E.1 Space Definitions and Mapping Framework
-We define four key spaces for formalizaing our alignment framework:
-  • Attribute space $\mathcal{B}\subset\mathbb{R}^{d}$: each $b\in\mathcal{B}$ is a structured configuration of interpretable attributes drawn from the attribute bank (e.g., ideology, altruism, personality).
-Attributes serve as control handles that steer LLM behavior.
-  • Endowment space $\mathcal{E}$: each $e\in\mathcal{E}$ is a natural language prompt or agent description instantiated from an attribute configuration $b\in\mathcal{B}$.
-Endowments serve as the actual inputs to the LLM.
-We treat $\mathcal{E}$ as a structured set, not a Euclidean space.
-  • Preference space $\mathcal{P}\subset\mathbb{R}^{m}$: each $p\in\mathcal{P}$ represents a latent vector encoding a behavioral inclination or preference profile.
-This space is unobservable but assumed to mediate how attributes shape responses.
-  • Response space $\mathcal{R}\subset\mathbb{R}^{n}$: each $r\in\mathcal{R}$ is an observable vector of responses to survey questions (e.g., binary or Likert-coded).
-Based on the four spaces, we define three mappings:
-  • Attribute-to-Endowment Mapping ($\phi:\mathcal{B}\to\mathcal{E}$): This is a stochastic mapping that transforms a symbolic attribute configuration into a textual prompt or persona instantiation readily parsable by an LLM.
-Since multiple distinct prompts can be generated from the same attributes, we model $\phi$ as sampling from a conditional distribution over prompts given attributes.
-  • Endowment-to-Preference Mapping ($g:\mathcal{E}\to\mathcal{P}$): This captures the internal LLM process by which a given prompt (endowment) induces a latent preference state.
-This mapping is governed by the LLM’s internal representations, shaped jointly by its large-scale pretraining corpora, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF).
-Albeit unobservable, this preference vector mediates the model’s behavioral tendencies in downstream response generation.
-  • Preference-to-Response Mapping ($h:\mathcal{P}\to\mathcal{R}$): This final mapping generates a survey response vector based on the latent preference state.
-It represents how an agent’s inclinations are expressed as answers to survey questions.
-E.2 Preference Reconstruction via a Functional Basis
-The overall agent behavior function is thus a composition of the three mappings:
-$f=h\circ g\circ\phi:\mathcal{B}\rightarrow\mathcal{E}\rightarrow\mathcal{P}\rightarrow\mathcal{R}$
-where $f(b)$ denotes the response behavior of an LLM agent endowed based on attributes $b$.
-Given a set of sampled agents $\mathcal{A}=\{b_{1},\dots,b_{k}\}$, we obtain a response basis $\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}$. We then treat alignment as a projection task:
-$\min_{w\in\mathbb{R}^{|\mathcal{A}|}}\left\|\sum_{b\in\mathcal{A}}w_{b}f(b)-r_{\text{pop}}\right\|^{2}$ (8)
-where $r_{\text{pop}}$ is the target aggregate response vector, and the weights $w_{b}$ reflect the inferred contribution of each proxy agent to the population preference.
-E.3 Alignment as a Two-Stage Problem
-We frame the alignment process as a two-stage problem grounded in mappings across attribute, endowment, preference and responses spaces.
-Stage 1: Constructing a functional basis based on response variability.
-The goal of the first stage is to construct a diverse and expressive set of LLM agents whose latent preferences collectively span the space where the ground-truth population preference resides.
-Since latent preferences are not directly observable, we use response variability as an indirect gauge, leveraging the preference-to-response mapping ($h:\mathcal{P}\to\mathcal{R}$).
-Starting from a sampled set of attribute configurations $\mathcal{A}=\{b_{1},\dots,b_{k}\}\subset\mathcal{B}$, we apply the composition mapping $f:\mathcal{B}\to\mathcal{R}$ to obtain a set of simulated agent responses:
-$\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}\subset\mathcal{R}.$
-We interpret variability in $\hat{\mathcal{R}}$ as an indirect signal of the latent variation in $\hat{\mathcal{P}}$.
-The design objective is to ensure that the span of $\hat{\mathcal{R}}$ is sufficiently rich so that the latent preference vectors form a functional basis for reconstructing population-level behavior.
-To optimize this objective, we operate in the more interpretable and tractable attribute space $\mathcal{B}$ rather than the high-dimensional preference space $\mathcal{P}$.
-By actively exploring semantically meaningful regions of $\mathcal{B}$—analogous to domain-aware active learning strategies Hao et al.
-[2024] and latent-manifold transformations Jin et al.
-[2024]—we adaptively improve coverage of $\hat{\mathcal{P}}$.
-Stage 2: Projecting onto population responses.
-In the second stage, we assume access to a population-level response vector $r_{\text{pop}}\in\mathcal{R}$, such as aggregate survey results.
-The construction of the agent ensemble is done by solving the optimization problem defined by (8), where each weight $w_{b}$ reflects the inferred contribution of the agent defined by attribute set $b$ in reconstructing the aggregate signal.
-Together, these two stages—constructing a coverage-maximizing basis and estimating population-aligned weights—enable efficient and demographically-agnostic preference reconstruction using steerable LLM agents.
+Appendix F Limitations and Future Work
+Question weighting and regression loss
+By design, P2P converts all multiple choice questions into binary questions before fitting a regression model.
+For example, a question with five options will turn into five questions with binary options.
+While this design ensures that regression can run smoothly with heterogeneous question types, it steers the model to prioritize fitting questions with more answer options, e.g., a 10-point Likert scale.
+Put differently, information contained in questions with fewer answer options, e.g., Yes/No, gets diluted after the binarization.
+Whether or not this design is desirable depends on the user’s assumption of the relationship between preference signal and choice granularity.
+The current design implicitly posits a positive relationship between the two—i.e., the finer the scale, the better a signal on the latent preference.
+For other assumptions, question reweighting or format transformation need to be done before the regression step.
+P2P uses option-level weighting as the default, but also offers block weighting as an option.
+Survey data in, text out
+The current design of P2P limits its use case to labeling: Given question and options, the aligned agent ensemble expresses its preference by offering a probability distribution over the given options.
+A key future research direction is to extend the use case to text completion, e.g., drafting resolution, offering suggestion and answering freeform survey questions.
+How to conflate individual agent textual responses into a representative aggregate response remains an under-explored area of research.
+Auction design (cf.
+Dütting et al., 2024) can be a promising aggregation algorithm in this case.
+P2P as a benchmark for model steerability
+In this study our primary goal is understanding preference reconstruction theory and its implementation through the two-stage alignment framework.
+While our focus is on the design aspects of the system, we acknowledge that the backend model’s steerability is also a determinant of P2P’s empirical performance.
+The ablation study using different backends offers first insights on this aspect.
+It also indicates P2P’s potential of being transformed into a benchmark pipeline for the study of model steerability in pluralistic alignment.
+Temporal Alignment and Non-stationary Preference
+Any method trained at a single moment risks becoming stale as attitudes and beliefs shift.
+This also affects P2P when considering using a fitted ensemble in future surveys.
+Our view is that synthetic agents should be periodically re-anchored using more recent, easier-to-collect data (e.g., shorter or cheaper surveys) and then tested on more demanding items.
+Architecturally, P2P is set up for this.
+The endowment generation and regression stages can be rerun on new waves, and one can track how learned weights and entropy patterns evolve, or when reconstruction quality begins to deteriorate.
+While P2P provides the machinery, systematic studies of preference shift and alignment across time and topics remain open.
+Short-cut Learning and Generalization
+Overfitting to survey artifacts is a key concern for any survey-based evaluation.
+This concern is closely related to what the robustness literature calls shortcut learning or Clever Hans behavior, where models exploit superficial regularities of a dataset instead of learning the underlying concept (Geirhos et al., 2020; Lapuschkin et al., 2019).
+Our empirical evaluation across 14 ATP waves is designed to mitigate the risk that our findings hinge on idiosyncrasies of a single wave, and the additional WVS experiments further mitigate the risk of ATP-specific shortcutting.
+While these steps serve as initial evidence against program-specific shortcutting, it does not constitute a complete solution to the generalization problem.
+A full multi-survey study (e.g., across GlobalOpinionQA and other polling corpora) is beyond the scope of this paper, but P2P ’s modular pipeline readily supports this kind of analysis.
+Fundamentally, we see this as partly a survey design challenge: if we want synthetic agents to learn stable preferences rather than survey-specific shortcuts, future work may need survey instruments that are explicitly constructed with learnability and cross-topic generalization in mind.
+In the last paragraph of Appendix , we suggest a semantic question module for P2P to quantify the expressiveness of survey questions, which could serve as a parallel line of work to mitigate shortcutting on the model-design side.
+Taken together, these directions highlight an important and promising interdisciplinary research frontier in our pursuit of alignment science.